### v1

# JEL classification: C33, C38, C55

This paper proposes a novel method for sparse latent factor modeling using a new sparse asymptotic Principal Component Analysis (APCA). This approach analyzes the co-movements of large-dimensional panel data systems over time horizons within a general approximate factor model framework. Unlike existing sparse factor modeling approaches based on sparse PCA, which assume sparse loading matrices, our sparse APCA assumes that factor processes are sparse over the time horizon, while the corresponding loading matrices are not necessarily sparse. This development is motivated by the observation that the assumption of sparse loadings may not be appropriate for financial returns, where exposure to market factors is generally universal and non-sparse. We propose a truncated power method to estimate the first sparse factor process and a sequential deflation method for multi-factor cases. Additionally, we develop a data-driven approach to identify the sparsity of risk factors over the time horizon using a novel cross-sectional cross-validation method. Theoretically, we establish that our estimators are consistent under mild conditions. Monte Carlo simulations demonstrate that the proposed method performs well in finite samples. Empirically, we analyze daily stock returns for a balanced panel of S&P 500 stocks from January 2004 to December 2016. Through textual analysis, we examine specific events associated with the identified sparse factors that systematically influence the stock market. Our approach offers a new pathway for economists to study and understand the systematic risks of economic and financial systems over time.
Keywords: Factor Analysis, Asymptotic Principal Components, Sparsity, Large-Dimension, Panel Data
JEL classification: C33, C38, C55
1 Introduction
In the realm of big-data analysis, the study of large-dimensional panel data has gained significant prominence in recent years. Large-dimensional panel data refers to datasets where observations are recorded for multiple individuals or entities over time, resulting in a wealth of information over the spatial and time horizons. Analyzing such datasets is essential for understanding complex economic, financial, and social phenomena. However, as the dimensionality of the data increases, traditional statistical techniques encounter numerous challenges, including multicollinearity, computational complexity, and difficulties in extracting meaningful insights. To address these challenges, latent factor modeling, also referred to as statistical factor modeling in Campbell, Lo, and MacKinlay (1997), has emerged as a powerful and versatile approach in the field of panel data analysis. Latent factor models offer a structured framework for dimension reduction and capturing common sources of variation in high-dimensional data. For example, asset returns in finance are often modeled as functions of a small number of factors; see Ross (1976) and Connor and Korajczyk (1986, 1988). Macroeconomic variables of multiple countries are found to have common components; see Stock and Watson (1989), Gregory and Head (1999), and Forni, Hallin, Lippi, Reichlin, et al. (2000). In demand systems, the Engle curves can be expressed in terms of a finite number of factors; see Lewbel (1991). As the dimensions of the panel data systems increase, various factor models have been developed to reduce the dimensionalities under different scenarios. See, for example, the approximate factor models in Chamberlain and Rothschild (1983), Stock and Watson (2002b, a), Bai and Ng (2002), Bai (2003), Fan, Liao, and Mincheva (2013), Lettau and Pelger (2020), Pelger (2020), and Gao and Tsay (2023, 2024), and the dynamic factor models in Forni, Hallin, Lippi, and Reichlin (2000), among others. These models are particularly well-suited for large-dimensional panel datasets where the number of variables (or features) can be moderately large.
One of the most profound challenges in factor modeling is to handle high-dimensional time series data efficiently and effectively. Sparse factor modeling represents a groundbreaking approach to this challenge, offering a powerful technique to extract meaningful insights from complex datasets characterized by a multitude of variables. Sparse factor modeling seeks to identify and capture the underlying structure of data while simultaneously promoting simplicity by selecting only a subset of relevant variables (or features) from the original set. The central idea behind sparse factor modeling is to uncover latent factors that drive the observed data’s variation while enforcing a degree of sparsity, meaning that only a few variables are deemed essential to explain the data’s structure. This approach is particularly valuable in scenarios where the number of variables far exceeds the number of observations, as it not only reduces computational demands but also enhances the interpretability of the results. Examples of articles concerning this approach include the sparse PCA in Jolliffe, Trendafilov, and Uddin (2003), Zou, Hastie, and Tibshirani (2006), Shen and Huang (2008), and Witten, Tibshirani, and Hastie (2009), and the sparse factor analysis with sparse loadings in Kristensen (2017), among others, where all of the aforementioned works are in line with factor analysis with sparse loadings, implying that each factor process is a linear combination of a small subset of the original panel series only.
It is widely known that the key idea in a large-dimensional factor model is that the dimensions of both the timeline and the cross-section of the data are large, and most of the co-movements can be explained by a few factors. These factors and loadings are usually estimated by the conventional principal component analysis (PCA) method due to its ability to parsimoniously capture much of the information in a large number of variables. However, the resulting latent PCs or factors are typically linear combinations of of all cross-sectional units/variables, which are usually hard to interpret. The sparse PCA restricts the cardinality of the weight vectors for the PCs, or equivalently, the loadings for the factor processes, so the the PCs are sparse linear combinations of the underlying variables. Consequently, the PCs or factors are only linear combinations of a small subset of the cross-sectional units, which facilitates the interpretations of the resulting PCs or factors. Nevertheless, there is little literature concerning the justificaton of the sparsity assumption in the loadings. For example, the assumption may not be appropriate for financial returns, where the exposure of the returns to a market factor is universal and non-sparse, as pointed out in Pelger and Xiong (2022), see also the empirical evidence in Gao and Tsay (2021).
In view of the above discussion, this paper marks a further development in the sparse factor modeling of large-dimensional panel data from a different perspective. Motivated by the empirical success of a general approximate factor model, see Pelger (2020) and the references therein, we also adopt such a similar framework. Unlike the sparse factor modeling approach with sparse loadings in existing literature, we assume the latent factor processes are sparse over time, while the loadings can be nonzero in general settings. Specifically, for a given time period of economic or financial series, we postulate that cross-sectional units may exhibit varying strengths of co-movement over time. During certain periods, the co-movement may not be significant enough, allowing idiosyncratic terms to dominate the co-movement, potentially resulting in the absence of a systematic risk factor. See the empirical evidence of the dominance of idiosyncratic risk in Campbell, Lettau, Malkiel, and Xu (2001) and Goyal and Santa-Clara (2003), among others. Consequently, the loading vector of each panel series can be treated as a linear combination of the corresponding series over the entire timeline. When a factor process is sparse, the corresponding loading associated with each series is a linear combination of the series over a small subset of the entire time horizon. This implies that only specific data points contribute to the strength of the loading for each series. The interpretation of this approach is quite straightforward. For example, asset returns in finance may not always exhibit systematic co-movements over the entire timeline, even though the market beta associated with the market factor can be significantly different from zero since it is calculated based on the entire timeline. However, clear and significant co-movement patterns do emerge during certain periods, such as when the Federal Reserve raises interest rates to combat inflation or during financial crises triggered by currency issues, debt, banking problems, or subprime mortgage defaults. Therefore, our proposed approach can facilitate researchers in economics and finance in discovering the connections between specific policies or events and the co-movements of economic or financial systems.
In this paper, we first introduce approximate factor models and their estimation procedures using the conventional PCA or the asymptotic PCA (APCA) method introduced by Connor and Korajczyk (1986, 1988). We then present a sparse APCA technique by formulating a sparse factor modeling framework with sparse factors over time horizons. Under the $\ell_{0}$-constraint imposed on the factors, we propose a truncated power method to estimate the sparse factors in the one-factor case, corresponding to the hard-thresholding technique in the machine learning literature. Additionally, we develop a sequential deflation estimation procedure for multi-factor cases, utilizing the truncated power method along with certain projection techniques. The theoretical properties of the proposed estimators are established under some mild conditions when both the dimension and the sample size go to infinity. Simulated and real data examples are used to illustrate the proposed method. To the best of our knowledge, this is the first study in the literature to consider sparse factors over the timeline, and we establish theoretical guarantees for the estimation procedure. The proposed approach provides an effective way to bridge time and significant risk factors, which play a crucial role in studying the dynamics of large-dimensional panel systems.
In an empirical application, we apply the proposed method to study the dynamics of daily stock returns for a balanced panel of S&P 500 stocks from January 2004 to December 2016. Through textual analysis, we examine specific events associated with the identified sparse factors that systematically influence the stock market. Our findings reveal nine significant time factors that may systematically affect the financial market: Economic Indicators, Government Policies, Global Events, Market Sentiment, Company-specific Factors, Credit Risk, Oil Price, China, and Europe. The findings contribute to a deeper understanding of market dynamics and the critical events that drive stock market movements, providing a valuable resource for both investors and policymakers in navigating the complexities of financial markets. The empirical study suggests that our proposed method not only provides interpretable results but also establishes connections between co-movements and specific events over time, effectively bridging time and risk. Our approach offers a new pathway for economists to study and understand the systematic risks of economic and financial systems over time.
Note that the proposed framework not only marks a clear distinction from existing sparse PCA or sparse factor modeling in the literature in terms of formulation and interpretation, but the technique used in this paper also differs from those employed in previous works. For example, in the sparse PCA literature, such as Zou, Hastie, and Tibshirani (2006), Shen and Huang (2008), and Witten, Tibshirani, and Hastie (2009), the true sparse eigenvectors are assumed to be constant vectors, and regularized estimation algorithms are proposed to estimate these sparse eigenvectors. In the realm of sparse factor modeling, such as in Kristensen (2017) and Uematsu and Yamagata (2022), the loadings are assumed to be sparse and non-random, with $\ell_{1}$-regularized methods used to estimate the sparse loadings. Conversely, in the sparse proximate factor modeling approach of Pelger and Xiong (2022), only the factor weights are assumed to be sparse, while the resulting loading and factor processes are non-sparse. In contrast with the aforementioned frameworks, the proposed sparse APCA method assumes that the random factor processes exhibit a sparse structure, meaning each factor data point can be either a random variable or zero over the timeline. A zero factor point indicates that the return series may not have co-movement at that specific time point. Our estimation procedure is developed under $\ell_{0}$-constraints on the factors. Therefore, our theoretical analysis differs from all the previously mentioned approaches, as the dependence structure along the timeline differs from that across the cross sections. Consequently, our method offers a way to link specific time points with the relevant risk factors that systematically influence the panel of interest.
This paper is also related to the literature on high-dimensional factor models with weak factors. In this context, there are two classes of weak factors: the first type affects only a subset of cross-sectional units, while the second type has relatively weak effects on the entire cross-section. The former type is often estimated using a group structure or some prior knowledge about factor strength, as discussed in Bailey, Kapetanios, and Pesaran (2021) and Anatolyev and Mikusheva (2022). In addition, Huang, Jiang, Li, Tong, and Zhou (2022) proposed a scaled PCA method to scale down weak factors through a supervised learning approach. This method was later extended to a supervised dynamic PCA method by Gao and Tsay (2024) for dynamic forecasting with weak factors. Importantly, Bai and Ng (2023) showed that PCA estimators exhibit a slower convergence rate under a weak factor model. Our paper fits into the second category, where factors over the time horizon are assumed to be either zero or nonzero, denoting weak and strong factors, respectively. Zero factors indicate that idiosyncratic components are relatively more dominant compared to weak systematic risk factors. For long time series sequences, our method also offers a dimension-reduction approach (in time horizon) to studying the dynamics of the economic and financial series over time horizons.
The contributions of this paper are multi-fold. First, to the best of our knowledge, this is the first study to consider sparse factors over the time horizon, providing economists with a novel approach to bridging the gap between time and risk in economic and financial systems. Second, unlike the $\ell_{1}$-penalized methods in sparse PCA, we propose using a truncated power method based on an $\ell_{0}$-constraint for the one-factor case and a sequential optimization procedure for estimating multiple factors. This extends the asymptotic PCA technique developed by Connor and Korajczyk (1986, 1988) for factor analysis to high dimensions. Third, rather than using the traditional cross-validation method to identify sparsity structure, we introduce a new cross-validation approach by partitioning the spatial dimensions, enabling the estimation of sparsity parameters through this method. Theoretically, we establish the consistency of the proposed estimators as the dimension and sample size approach infinity. Importantly, we provide theoretical guarantees for the proposed cross-sectional cross-validation method, demonstrating its ability to consistently estimate the sparsity structure over the time horizon of factor models. This offers a rigorous approach to linking time and significant systematic risks across the time horizon.
The rest of the paper is organized as follows. Section 2 introduces the sparse factor modeling framework and its estimation procedure. Section 3 presents asymptotic properties of the estimators obtained in Section 2. Section 4 studies the finite-sample performance of the proposed approach via simulation, and 5 illustrates the proposed procedure with an empirical application. Section 6 concludes. All the proofs and derivations for the asymptotic results are relegated to an online Appendix.
Notation: We use the following notation. For a $p\times 1$ vector ${\mathbf{u}}=(u_{1},...,u_{p})^{\prime}$, $\|{\mathbf{u}}\|_{1}=\sum_{i=1}^{p}|u_{i}|$ is the $\ell_{1}$-norm and $\|{\mathbf{u}}\|_{\infty}=\max_{1\leq i\leq p}|u_{i}|$ is the $\ell_{\infty}$-norm. ${\mathbf{I}}_{p}$ denotes the $p\times p$ identity matrix. For a matrix ${\mathbf{H}}$, its Frobenius norm is $\|{\mathbf{H}}\|=[\mathrm{trace}({\mathbf{H}}^{\prime}{\mathbf{H}})]^{1/2}$ and its operator norm is $\|{\mathbf{H}}\|_{2}=\sqrt{\lambda_{\max}({\mathbf{H}}^{\prime}{\mathbf{H}})}$, where $\lambda_{\max}(\cdot)$ denotes the largest eigenvalue of a matrix, and $\|{\mathbf{H}}\|_{\min}$ is the square root of the minimum non-zero eigenvalue of ${\mathbf{H}}{\mathbf{H}}^{\prime}$. $|{\mathbf{H}}|$ denotes the absolute value of ${\mathbf{H}}$ elementwisely. The superscript denotes the transpose of a vector or matrix. We also use the notation $a\asymp b$ to denote $a=O(b)$ and $b=O(a)$ or $a$ and $b$ have the same order of stochastic bound when they are random variables.
2 Model and Methodology
2.1 Model Setup
Let $x_{i,t}$ be the $i$-th unit of the cross-sectional panel ${\mathbf{x}}_{t}=(x_{1,t},...,x_{N,t})^{\prime}\in R^{N}$ at time $t$, for example, $x_{i,t}$ can be the stock return of the $i$-th asset at time $t$, we consider the following approximate factor model:
$x_{i,t}=\boldsymbol{\lambda}_{i}^{\prime}{\mathbf{f}}_{t}+e_{i,t},$ (1)
where $x_{i,t}$ is the only observed datum for the $i$-th cross-section at time $t$ ($i=1,...,N$; $t=1,...,T$), ${\mathbf{f}}_{t}$ is an $r$-dimensional vector of common or systematic risk factors, $\boldsymbol{\lambda}_{i}$ is an $r$-dimensional vector of factor loadings, and $e_{i,t}$ is the idiosyncratic component of $x_{i,t}$ that cannot be explained by the common risk factors. We may combine all the cross-sectional units together and write the above equation as
${\mathbf{x}}_{t}=\boldsymbol{\Lambda}{\mathbf{f}}_{t}+{\mathbf{e}}_{t},$ (2)
where $\boldsymbol{\Lambda}=(\boldsymbol{\lambda}_{1},...,\boldsymbol{\lambda}_{N})^{\prime}$ and ${\mathbf{e}}_{t}=(e_{1,t},...,e_{N,t})^{\prime}$. Assume a panel data set of $T$ time-series observations and $N$ cross-sectional observations, denoted as ${\mathbf{X}}\in\mathbb{R}^{T\times N}$, has a factor structure with $r$ common factors. Let ${\mathbf{F}}=({\mathbf{f}}_{1},...,{\mathbf{f}}_{T})^{\prime}$ and ${\mathbf{e}}=({\mathbf{e}}_{1},...,{\mathbf{e}}_{T})^{\prime}$, then Models (1) and (2) can be written in the following matrix form
${\mathbf{X}}={\mathbf{F}}\boldsymbol{\Lambda}^{\prime}+{\mathbf{e}}.$ (3)
For ease of notation, we also denote ${\mathbf{X}}=(\underline{{\mathbf{x}}}_{1},...,\underline{{\mathbf{x}}}_{N})$, ${\mathbf{F}}=(\underline{{\mathbf{f}}}_{1},...,\underline{{\mathbf{f}}}_{r})$, and $\boldsymbol{\Lambda}=(\underline{\boldsymbol{\lambda}}_{1},...,\underline{\boldsymbol{\lambda}}_{r})$ when referring to their columns. In the econometrics/statistical and finance literature, Model (3) is usually estimated using the Principal Components or Asymptotic Principal Components estimation method (see Connor and Korajczyk (1986, 1988), Bai and Ng (2002), Bai (2003), and Fan, Liao, and Mincheva (2013), among others). As a result, all the elements in the estimators $\widehat{\mathbf{F}}$ and $\widehat{\boldsymbol{\Lambda}}$ are usually nonzero, making it difficult to interpret the dynamic and cross-sectional relationships of the data. Moreover, the role of the idiosyncratic terms is rarely considered or even ignored in characterizing the individual dynamics of the series.
In the next subsection, we briefly introduce the identification issues in Model (3) and the estimators obtained by the method of PCA.
2.2 Identification and PCA Estimation
As the panel of data ${\mathbf{X}}$ is the only observable component in Model (3), and ${\mathbf{F}}$ and $\boldsymbol{\Lambda}$ are latent and unobserved ones, the pair $({\mathbf{F}},\boldsymbol{\Lambda})$ itself is not identifiable since they can be replaced by $({\mathbf{F}}{\mathbf{H}}^{\prime},\boldsymbol{\Lambda}{\mathbf{H}}^{-1})$ for any $r\times r$ non-singular matrix ${\mathbf{H}}$, without altering the equation in (3). Therefore, it is conventional to impose some constraints on the factors and their associated loadings. There are two equivalent sets of identification conditions in the PCA framework. The first one assumes
$\textnormal{Cov}({\mathbf{f}}_{t})={\mathbf{I}}_{r}\,\,\text{and}\,\,\boldsymbol{\Lambda}^{\prime}\boldsymbol{\Lambda}\,\,\text{is diagonal},$ (4)
and the second one postulates that
$\frac{1}{N}\boldsymbol{\Lambda}^{\prime}\boldsymbol{\Lambda}={\mathbf{I}}_{r}\,\,\text{and}\,\,\textnormal{Cov}({\mathbf{f}}_{t})\,\,\text{is diagonal}.$ (5)
See, for example, Bai and Ng (2002) and Fan, Liao, and Mincheva (2013). Note that ${\mathbf{F}}$ and $\boldsymbol{\Lambda}$ are still not identifiable when the diagonals in (4) or (5) have multiplicity or there exists a diagonal matrix ${\mathbf{H}}$ with the diagonals being $1$ or $-1$ such that $({\mathbf{F}}{\mathbf{H}},\boldsymbol{\Lambda}{\mathbf{H}})$ is still a pair of solutions. In this paper, we do not consider the special case of multiplicities, and the change of signs does not affect our analysis because it gives us an advantage to choose an ${\mathbf{F}}$ with any particular signs to facilitate the interpretations of the estimates. On the other hand, according to the discussion in Bai and Ng (2013), the estimates can be uniquely defined asymptotically under conditions (4) or (5) if the top eigenvalues of the covariance of the panel are distinct. They showed that the rotation matrix ${\mathbf{H}}$ is asymptotically equal to an identity with certain convergence rates. As a matter of fact, the conditions in (4) and (5) are reasonable if we adopt the PCA method to estimate the factors and their associated loadings, because the estimated PCA factors $\widehat{\mathbf{F}}$ and loading $\widehat{\boldsymbol{\Lambda}}$ either satisfy the condition in (4) or (5) depending on whether we apply PCA to ${\mathbf{X}}{\mathbf{X}}^{\prime}$ or ${\mathbf{X}}^{\prime}{\mathbf{X}}$ as discussed below.
Suppose ${\mathbf{x}}_{t}$ is a demeaned vector, under condition (4), the estimation of ${\mathbf{F}}$ is obtained by conducting a PCA on ${\mathbf{X}}{\mathbf{X}}^{\prime}$ or a scaled version $\frac{{\mathbf{X}}{\mathbf{X}}^{\prime}}{NT}$, and each estimated loading vector $\boldsymbol{\lambda}_{i}$ is a linear combination of the $i$th series $\underline{{\mathbf{x}}}_{i}=(x_{i,1},...,x_{i,T})^{\prime}$ using the weight matrix $\widehat{\mathbf{F}}$. Specifically, under condition (4), the estimated factor matrix, denoted by $\widehat{\mathbf{F}}$, is $\sqrt{T}$ times the normalized eigenvectors corresponding to the $r$ largest eigenvalues of the $T\times T$ matrix ${\mathbf{X}}{\mathbf{X}}^{\prime}$. For a given $\widehat{\mathbf{F}}$, the estimated factor loading matrix is $\widehat{\boldsymbol{\Lambda}}^{\prime}=\widehat{\mathbf{F}}^{\prime}{\mathbf{X}}/T$, which can be obtained by ordinary least-squares (OLS) estimation. This approach is also known as Asymptotic Principal Component Analysis (APCA) in the finance literature; see Connor and Korajczyk (1986, 1988) for details. Consequently, the estimated loading vector $\boldsymbol{\lambda}_{i}$ for the $i$-th series is a linear combination of the $i$-th series $\underline{{\mathbf{x}}}_{i}=(x_{i,1},...,x_{i,T})^{\prime}$ using the weight matrix $\widehat{\mathbf{F}}$ over the timeline.
Similarly, under condition (5), the estimation of $\boldsymbol{\Lambda}$ is obtained by performing a PCA on ${\mathbf{X}}^{\prime}{\mathbf{X}}$ (or $\frac{{\mathbf{X}}^{\prime}{\mathbf{X}}}{NT}$). By abuse of notation, we still denote the estimated loading and factor matrices as $\widehat{\boldsymbol{\Lambda}}$ and $\widehat{\mathbf{F}}$, respectively. The estimated loading matrix $\widehat{\boldsymbol{\Lambda}}$ is constructed as $\sqrt{N}$ times the normalized eigenvectors corresponding to the $r$ largest eigenvalues of the $N\times N$ matrix ${\mathbf{X}}^{\prime}{\mathbf{X}}$. Then, the estimated factor matrix is $\widehat{\mathbf{F}}={\mathbf{X}}\widehat{\boldsymbol{\Lambda}}/N$. Therefore, each factor $\widehat{\mathbf{f}}_{t}$ is a linear combination of ${\mathbf{x}}_{t}=(x_{1,t},...,x_{N,t})$ using the weights in $\widehat{\boldsymbol{\Lambda}}$ over the cross section.
We highlight that the conditions in (4) and (5) are only imposed in order to estimate the factors and loading matrix by the PCA method, and the true underlying factors and loading matrix do not necessarily satisfy either (4) or (5) so long as some regularity conditions as those in Assumptions A-B of Bai and Ng (2002) are met. For example, if the estimated factors satisfy that $\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}}/T={\mathbf{I}}_{r}$, then there exists an invertible rotation matrix ${\mathbf{H}}\in R^{r\times r}$ such that the consistency of $\widehat{\mathbf{F}}$ to ${\mathbf{F}}{\mathbf{H}}^{\prime}$ can be established; See Bai (2003) and Fan, Liao, and Mincheva (2013) for details. Moreover, if we further assume the true factors satisfy ${\mathbf{F}}^{\prime}{\mathbf{F}}/T={\mathbf{I}}_{r}$, or more rigorously, $\textnormal{Cov}({\mathbf{f}}_{t})={\mathbf{I}}_{r}$ as that in (4), then the rotation matrix ${\mathbf{H}}$ is an orthogonal one (asymptotically) as shown in Fan, Liao, and Mincheva (2013). If we additionally assume the loading matrix satisfy the condition in (4), then ${\mathbf{H}}$ is an identity asymptotically, as shown in Bai and Ng (2013).
Note that the two estimation procedures are equivalent in the sense that the estimated common component $\widehat{\mathbf{F}}\widehat{\boldsymbol{\Lambda}}^{\prime}$ stays the same, which is due to the nature of the PCA method. In either case, the PCA factors and loadings are difficult to interpret. This poses a challenge for modeling and understanding the dynamic structure and the co-movement in the data, especially for those in finance and economics where understanding the hidden economic mechanism in the data is a main objective to study.
In this article, we adopt the conditions in (4) and pick a specific choice of the rotation matrix based on the APCA approach, ensuring the unique identification of ${\mathbf{F}}$ and $\boldsymbol{\Lambda}$. For example, suppose the true underlying factor and loading matrices are ${\mathbf{F}}^{*}$ and $\boldsymbol{\Lambda}^{*}$, respectively, such that ${\mathbf{X}}={\mathbf{F}}^{*}\boldsymbol{\Lambda}^{*}{{}^{\prime}}+{\mathbf{e}}$, we can show that there exists a rotation matrix ${\mathbf{H}}$ such that the estimated APCA factors $\widehat{\mathbf{F}}$ with $\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}}/T={\mathbf{I}}_{r}$ is consistent to ${\mathbf{F}}^{*}{\mathbf{H}}^{\prime}$ (up to a scale). Hence, by redefining ${\mathbf{F}}={\mathbf{F}}^{*}{\mathbf{H}}^{\prime}$ and $\boldsymbol{\Lambda}=\boldsymbol{\Lambda}^{*}{\mathbf{H}}^{-1}$, the model transforms into ${\mathbf{X}}={\mathbf{F}}\boldsymbol{\Lambda}^{\prime}+{\mathbf{e}}$, enabling us to uniquely estimate ${\mathbf{F}}$ and $\boldsymbol{\Lambda}$ through the APCA approach. With this redefined framework, our sparsity assumption will be directly imposed on ${\mathbf{F}}$ in subsequent analysis, as there are no identification issues as discussed above. Our work aims to extend the APCA of Connor and Korajczyk (1986, 1988) to high dimensions using advanced machine learning techniques.
2.3 Sparse Factor Processes
In view of the discussion in Section 2.2, the traditional APCA estimators are difficult to interpret, especially when the dimension is large. In the past decades, various sparse PCA or sparse factor modeling techniques have been proposed in the statistical and econometrics literature. See, for example, Zou, Hastie, and Tibshirani (2006), Witten, Tibshirani, and Hastie (2009), and Kristensen (2017), among other. Sparse PCA restricts the cardinality of the weight vectors for the PCs so that the PCs are sparse linear combinations of the underlying cross-sectional variables. By setting the factor weights to zero, sparse PCA facilitates the interpretation of the PCs, which improves interpretability without sacrificing the explanatory ability of the PCs or factors. To the best of our knowledge, all the aforementioned approaches assume that the loadings are sparse in the population model, and formulate the estimation of sparse loadings as a regularized optimization problem to estimate principle components with a penalty term, which allows for developing asymptotic inferential theory. Nevertheless, the assumption of sparse population loadings may not be satisfied in many datasets. For example, in finance, the exposure of asset returns to a market factor is universal and non-sparse. In addition, according to Camacho, Smilde, Saccenti, and Westerhuis (2020), the sparse PCA method presents disparate and limited performance when modeling sparse data, even in the optimistic case of noise-free data.
In this paper, unlike the existing approach in sparse factor modeling where the loadings are sparse, we assume the systematic risk factor process ${\mathbf{f}}_{t}$ is sparse along the time horizon as it characterizes the co-movement of the panel dynamically. When all components of ${\mathbf{f}}_{t}$ are zero at some certain time point, it implies that the panel is driven dominantly by individual factors or the corresponding idiosyncratic terms rather than a systematic one. On the other hand, for the economic and financial panel series, the nonzero factor ${\mathbf{f}}_{t}$ at certain data points can be treated as a systematic response of the panel to some important and influential economic or financial events or outcomes.
To illustrate the interpretations of the proposed approach, we take the panel of asset returns in finance as an example. First of all, the factor process measures the overall co-movement of all stock returns over the time horizon considered in the data set. When there is no co-movement over some certain period, the factors would be zero at such time points. Under such circumstances, the stock returns would be driven by their idiosyncratic terms because there are not enough government policies, influential news, or other major events that can drive all the cross-sectional movement, and consequently, the systematic response might be not significant. Second, each loading vector is a linear combination of the corresponding series over the time horizon, which measures the average dependence on the factors for the given time period of the data. When the factors are zero at a specific time point, the contribution of the series at such time point to the loading would be negligible.
2.4 Examples of Sparse Factors in Time Horizons
Factor analysis is an important statistical tool for reducing the dimensions of large panels of economic and financial data. When the factors are latent and unobservable, the conventional PCA estimation procedure, though easy to implement and possessing good theoretical properties (Bai and Ng (2002) and Bai (2003)), often results in loadings and factors that provide unclear economic insights and are difficult to interpret in practice. Under the proposed framework, sparse factors will improve the interpretability of the factors and offer a useful tool for understanding systematic risk over the timeline in economic and financial panel systems.
  1. (i)
    Asset Pricing Models. In the Arbitrage Pricing Theory (APT) proposed by Ross (1976), a fundamental assumption is that a small number of factors can explain a large number of asset returns. In contrast to cases where the factors are observable, as in Fama and French (1993), the latent factors ${\mathbf{f}}_{t}$’s extracted from the panel ${\mathbf{x}}_{t}$’s are difficult to interpret. Financial economists widely recognize that time and risk are two critical factors that make finance challenging (see, for example, Treynor (1961)). The finance subject would be incomplete without these two elements. The proposed framework provides an approach to bridge the systematic risk factor ${\mathbf{f}}_{t}$ and time $t$ for financial returns. When the systematic/common factors are nonzero at certain time points, we may relate events such as macroeconomic conditions, news, government policies, and banking conditions to the nonzero systematic risk factors detected over that time period. Conversely, when the systematic factors are zero, we may conclude that the market was mainly driven by company-specific shocks during those corresponding time periods.
  2. (ii)
    Disaggregate Business Cycle Analysis. Gregory and Head (1999) found that cross-country variations have common components, referred to as global shocks. In addition, each country also experiences country-specific shocks. It is clear that the global and country-specific shocks are reflected by the common factors ${\mathbf{f}}_{t}$’s and the idiosyncratic terms ${\mathbf{e}}_{t}$’s in Model (2), respectively. Under the proposed framework, sparse factors over a certain time period correspond to situations where global shocks do not play a dominant role in driving the cyclical variations in a country’s economy. Conversely, nonzero factors will inspire researchers to discover the possible reasons for significant global shocks during the corresponding time period.
  3. (iii)
    Consumption and Demand System. Let $x_{i,t}$ be the budget share of good $i$ for agent $t$, where we have $N$ goods and $T$ agents in total. Consumer theory assumes that $x_{i,t}=\boldsymbol{\lambda}_{i}^{\prime}{\mathbf{f}}(z_{t})+e_{i,t}$ in a single price regime, where ${\mathbf{f}}(z_{t})$ is an $r$-dimensional unknown functional factors of $z_{t}$ with $z_{t}$ being agent $t$’s total expenditure. See Section 4 of Lewbel (1991) for details. For certain agents, their budget shares for good $i$ can be zero, implying that ${\mathbf{f}}(z_{t})$’s are zero for $t$ belonging to such agents. Therefore, the proposed sparse factor models may provide economists with one way to understand certain groups of households’ preferences for specific goods. See a similar argument in Gabaix (2014).
  4. (iv)
    Monitoring and Forecasting. Stock and Watson (1998) and Forni, Hallin, Lippi, Reichlin, et al. (2000) have demonstrated that factor models provide an effective way to monitor economic activities, as large-dimensional economic panel data can be described by a small number of factors. In other words, business cycles can be modeled by the co-movements of economic variables, which correspond to the common factors ${\mathbf{f}}_{t}$’s in Model (2). When the factors are sparse, we can monitor economic activities according to the systematic risks detected by our approach. Additionally, Stock and Watson (1998) also considered forecasting inflation with the common factors (diffusion indices) extracted from a large panel of macroeconomic variables and found that these factors can improve forecasting accuracy.
2.5 Sparse APCA: Formulation and Estimation
We first introduce some further notation used in the estimation procedure. For an $N\times 1$ vector ${\mathbf{v}}=(v_{1},...,v_{N})^{\prime}$, $\|{\mathbf{v}}\|_{2}=\sqrt{\sum_{i=1}^{2}v_{i}^{2}}$ is the Euclidean norm and $\|{\mathbf{v}}\|_{0}=\text{card}\{\text{support}({\mathbf{v}})\}$ is the cardinality or number of non-zero elements in ${\mathbf{v}}$. Let $\mathbb{V}$ be the set of $N\times r$ orthogonal matrices and $\mathbb{V}^{\perp}$ be the set consisting of the orthogonal complements of the ones in $\mathbb{V}$. For a matrix ${\mathbf{V}}=({\mathbf{v}}_{1},...,{\mathbf{v}}_{r})\in\mathbb{V}$ and a vector ${\mathbf{s}}=(s_{1},...,s_{r})^{\prime}$, $\|{\mathbf{V}}\|_{0}\leq{\mathbf{s}}$ implies that $\|{\mathbf{v}}_{i}\|_{0}\leq s_{i}$ for $1\leq i\leq r$.
In the following subsection, we first formulate the estimation procedure in the one-factor case, i.e. the number of factors $r=1$ in Model (3), and the multi-factor case can be carried out based on the one-factor estimation procedure.
2.5.1 One-Factor Case
We consider the case when $r=1$ and the first factor process over the timeline is $\underline{{\mathbf{f}}}_{1}=(f_{1,1},...,f_{1,T})^{\prime}$. Let ${\mathbf{S}}={{\mathbf{X}}{\mathbf{X}}^{\prime}}/{(NT)}$, it follows that $\underline{{\mathbf{f}}}_{1}/\sqrt{T}$ can be approximated by the first normalized eigenvector of the $T\times T$ positive semi-definite matrix ${\mathbf{S}}$ according to Model (3). We assume $\|{\mathbf{v}}\|_{0}\leq s_{1}$, it suffices to solve the following optimization problem:
$\widehat{\mathbf{v}}=\arg\max_{{\mathbf{v}}^{\prime}{\mathbf{v}}=1}{\mathbf{v}}^{\prime}{\mathbf{S}}{\mathbf{v}},\,\,\text{subject to}\,\,\|{\mathbf{v}}\|_{0}\leq s_{1},$ (6)
and consequently, the estimator for $\underline{{\mathbf{f}}}_{1}$ is denoted as $\underline{\widehat{\mathbf{f}}}_{1}=\sqrt{T}\widehat{\mathbf{v}}$ such that $\underline{\widehat{\mathbf{f}}}_{1}^{\prime}\underline{\widehat{\mathbf{f}}}_{1}/T=1$ which satisfies the first condition in (4). Note that the solutions in (6) are the same as those obtained by PCA in Bai and Ng (2002) and Gao and Tsay (2023), among others, if the constraint $\|{\mathbf{v}}\|_{0}\leq s_{1}$ is removed. Furthermore, the problem in (6) is also equivalent to finding the largest eigenvalue associated with an $s_{1}$-sparse eigenvector of the matrix ${\mathbf{S}}$:
$\lambda_{\max}({\mathbf{S}},s_{1})=\max_{{\mathbf{v}}^{\prime}{\mathbf{v}}=1}{\mathbf{v}}^{\prime}{\mathbf{S}}{\mathbf{v}},\,\,\text{subject to}\,\,\|{\mathbf{v}}\|_{0}\leq s_{1},$ (7)
which is a non-convex problem in general. In fact, it is NP-hard because it can be reduced to the subset selection problem for ordinary least-squares problem; see Moghaddam, Weiss, and Avidan (2006).
In fact, the optimization in (6) is still a sparse PCA problem symbolically, and numerous methods have been developed to obtain sparse eigenvectors during the past decades. See the regularization method with elastic net in Zou, Hastie, and Tibshirani (2006) and the penalized matrix decomposition (PMD) algorithm using $\ell_{1}$-penalty in Witten, Tibshirani, and Hastie (2009), among others.
In this paper, we extend the APCA method of Connor and Korajczyk (1986, 1988) to high dimensions by adopting an $\ell_{0}$-constraint on the sparse factors. We propose to use the truncated power method introduced in Yuan and Zhang (2013) to solve the first normalized sparse eigenvector in problem (6). The method is similar to the classical power method but includes an additional truncation operation to ensure sparsity. The rationale for this is as follows. First, the conventional APCA is conducted based on the following matrix perturbation formulation:
${\mathbf{S}}=\boldsymbol{\Sigma}+{\mathbf{E}},$ (8)
where ${\mathbf{S}}$ is the matrix obtained by the noisy observation ${\mathbf{x}}_{t}$, $\boldsymbol{\Sigma}$ is a symmetric matrix whose eigenvectors are the true ones, and ${\mathbf{E}}$ is a random perturbation. The decomposition of (8) is formulated in (IA.1) of the online Appendix. If the true largest eigenvector ${\mathbf{v}}_{1}$ of $\boldsymbol{\Sigma}$ is sparse, then it is natural to recover ${\mathbf{v}}_{1}$ from the noisy matrix ${\mathbf{S}}$. This recovery is guaranteed when the error ${\mathbf{E}}$ is of a smaller order using the well-known $\sin\theta$ theorem in Davis and Kahan (1970) under the approximate factor model. Second, for any given vector ${\mathbf{u}}_{0}\in R^{N}$, the power method estimates the first eigenvector by
${\mathbf{u}}_{k}={\mathbf{S}}{\mathbf{u}}_{k-1}=...={\mathbf{S}}^{k}{\mathbf{u}}_{0},\,\,k=1,2,...,$ (9)
and a normalized ${\mathbf{u}}_{k}$ converges to the first eigenvector of ${\mathbf{S}}$. To see this, note that there exist $c_{1}$,…, $c_{N}$ such that
${\mathbf{u}}_{0}=c_{1}\widehat{\mathbf{v}}_{1}+c_{2}\widehat{\mathbf{v}}_{2}+...+c_{N}\widehat{\mathbf{v}}_{N},$
where $\widehat{\mathbf{v}}_{1},....,\widehat{\mathbf{v}}_{N}$ are the $N$ eigenvectors associated with the eigenvalues $\{\widehat{\lambda}_{i},i=1,...,N\}$ of ${\mathbf{S}}$. Then,
${\mathbf{u}}_{k}={\mathbf{S}}^{k}{\mathbf{u}}_{0}=\sum_{i=1}^{N}c_{i}{\mathbf{S}}^{k}\widehat{\mathbf{v}}_{i}=\widehat{\lambda}_{1}^{k}\left[c_{1}\widehat{\mathbf{v}}_{1}+\sum_{i=2}^{N}c_{i}(\frac{\widehat{\lambda}_{i}}{\widehat{\lambda}_{1}})^{k}\widehat{\mathbf{v}}_{i}\right].$
Under the assumption that $|\widehat{\lambda}|>|\widehat{\lambda}_{i}|$ for $i\geq 2$, it follows that the direction of ${\mathbf{u}}_{k}$ converges to that of $\widehat{\mathbf{v}}_{1}$.
Motivated by the above discussion, we modify the power method by adding a truncation in each iteration, and the procedure is given in Algorithm 1 below.
Algorithm 1 (Estimation procedure of the first sparse risk factor process).
1:Input: The scaled covariance matrix ${\mathbf{S}}={\mathbf{X}}{\mathbf{X}}^{\prime}/(NT)$, an initial vector ${\mathbf{u}}_{0}\in R^{T}$, a cardinality integer $s_{1}\in\{1,...,T\}$;
2:Let $t=1$;
3:repeat
4:     Compute ${\mathbf{u}}_{t}^{*}={\mathbf{S}}{\mathbf{u}}_{t-1}/\|{\mathbf{S}}{\mathbf{u}}_{t-1}\|$;
5:     Let $l_{t}=\text{supp}({\mathbf{u}}_{t}^{*},s_{1})$ be the indices of ${\mathbf{u}}_{t}^{*}$ with the largest $s_{1}$ absolute values;
6:     Compute $\widehat{{\mathbf{u}}}_{t}=\text{Truncate}({\mathbf{u}}_{t}^{*},l_{t})$, where it only keeps the elements in ${\mathbf{u}}_{t}^{*}$ with indices $l_{t}$;
7:     Normalize ${\mathbf{u}}_{t}=\widehat{{\mathbf{u}}}_{t}/\|\widehat{{\mathbf{u}}}_{t}\|$;
8:     $t\leftarrow t+1$;
9:until Convergence;
10:Output: $\widehat{\mathbf{v}}_{1}={\mathbf{u}}_{t}$.
The procedure, presented in Algorithm 1, generates a sequence of intermediate $s_{1}$-sparse eigenvectors $\{{\mathbf{u}}_{1},{\mathbf{u}}_{2},...\}$ from an initial sparse approximation ${\mathbf{u}}_{0}$. At each time stamp $t$, the intermediate vector ${\mathbf{u}}_{t-1}$ is multiplied by ${\mathbf{S}}$, and then the entries are truncated to zeros except for the largest $s_{1}$ entries. The resulting vector is then normalized to unit length. Finally, the estimated first factor process is $\underline{\widehat{\mathbf{f}}}_{1}=\sqrt{T}\widehat{\mathbf{v}}_{1}$, where $\widehat{\mathbf{v}}_{1}$ is the output one in Algorithm 1. Note that the sparsity parameter $s_{1}$ is unknown in practice, and we will propose a cross-validation procedure below to choose such a parameter.
In practice, the convergence criterion in Algorithm 1 needs to be given first. A common criterion for the convergence of Algorithm 1 is that the two eigenvector iterates ${\mathbf{u}}_{t}$ and ${\mathbf{u}}_{t-1}$ for $t\geq 1$ satisfy
$\|{\mathbf{u}}_{t}-{\mathbf{u}}_{t-1}\|_{\infty}\leq{{\varepsilon}},$ (10)
for a small ${{\varepsilon}}>0$. Simulation results suggest that the convergence is not sensitive to a sufficiently small ${{\varepsilon}}$, and the numerical results in Section 4 indicate that the algorithm works well when we take ${{\varepsilon}}=10^{-3}$ in (10).
2.5.2 Multi-Factor Case
In this section, we consider the general case when there are multiple factor processes in Model (3), i.e. $r>1$. In the presence of more than one common factor process, we need to apply Algorithm 1 multiple times and extract all the common factors in a sequential way. Specifically, when $\widehat{\mathbf{v}}_{1}$ is given via Algorithm 1, we subtract the projection on first factor component of the panel and the residual would be $\widetilde{\mathbf{X}}={\mathbf{X}}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{X}}=({\mathbf{I}}_{N}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}){\mathbf{X}}$. The resulting scaled covariance $\widetilde{\mathbf{S}}_{1}=({\mathbf{I}}_{N}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}){\mathbf{S}}({\mathbf{I}}_{N}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime})$. Then the second eigenvector $\widehat{\mathbf{v}}_{2}$ can be obtained by solving the following optimization problem (11) below:
$\widetilde{\mathbf{v}}_{2}=\arg\max{\mathbf{v}}^{\prime}\widetilde{\mathbf{S}}_{1}{\mathbf{v}},\,\,\text{subject to}\,\,{\mathbf{v}}^{\prime}({\mathbf{I}}_{N}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}){\mathbf{v}}=1\,\,\text{and}\,\,\|{\mathbf{v}}\|_{0}\leq s_{2},$ (11)
and the second eigenvector is $\widehat{\mathbf{v}}_{2}=\widetilde{\mathbf{v}}_{2}/\|\widetilde{\mathbf{v}}_{2}\|_{2}$. Consequently, the second estimated factor process is $\underline{\widehat{\mathbf{f}}}_{2}=\sqrt{T}\widehat{\mathbf{v}}_{2}$. We use the normalization condition ${\mathbf{v}}^{\prime}({\mathbf{I}}_{N}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}){\mathbf{v}}=1$ to maximize the additional variance of the original matrix ${\mathbf{S}}$, which is the same as the deflation framework in Mackey (2008). In other words, we need to eliminate the effect of the previous eigenvector directions by a projection method. Consequently, we formulate the procedure in Algorithm 2 below.
Algorithm 2 (A sequential estimation procedure of the sparse risk factors).
1:Input: The scaled covariance matrix ${\mathbf{S}}$, the cardinalities of $r$ columns $\{s_{1},...,s_{r}\}$ where $s_{i}\in\{1,2,...,T\}$;
2:Let $i=1$, ${\mathbf{B}}_{1}={\mathbf{I}}_{T}$;
3:repeat
4:     Given $s=s_{i}$, solve
$\widehat{\mathbf{v}}_{i}=\arg\min_{{\mathbf{v}}^{\prime}{\mathbf{B}}_{i}{\mathbf{v}}=1,\|{\mathbf{v}}\|_{0}\leq s_{i}}{\mathbf{v}}^{\prime}{\mathbf{S}}{\mathbf{v}},$
5:     Compute ${\mathbf{q}}_{i}={\mathbf{B}}_{i}\widehat{\mathbf{v}}_{i}$;
6:     Update ${\mathbf{S}}$ by ${\mathbf{S}}\leftarrow({\mathbf{I}}_{T}-{\mathbf{q}}_{i}{\mathbf{q}}_{i}^{\prime}){\mathbf{S}}({\mathbf{I}}_{T}-{\mathbf{q}}_{i}{\mathbf{q}}_{i}^{\prime})$;
7:     Update ${\mathbf{B}}_{i+1}={\mathbf{B}}_{i}({\mathbf{I}}_{T}-{\mathbf{q}}_{i}{\mathbf{q}}_{i}^{\prime})$.
8:     Return $\widehat{\mathbf{v}}_{i}\leftarrow\widehat{\mathbf{v}}_{i}/\|\widehat{\mathbf{v}}_{i}\|_{2}$.
9:     $i\leftarrow i+1$.
10:until $i=r+1$;
11:Output: $\{\widehat{\mathbf{v}}_{1},...,\widehat{\mathbf{v}}_{r}\}$;
From the optimization problem in (11) and the procedure in Algorithm 2, we cannot enforce the orthogonality and sparsity at the same time, which is similar to the case in the sparse PCA framework. When there is no $\ell_{0}$-constraint in (6) and (11), we can easily obtain that $\widehat{\mathbf{v}}_{1}$ and $\widehat{\mathbf{v}}_{2}$ are the two normalized and orthogonal eigenvectors associated with the top two eigenvalues of ${\mathbf{S}}$ and, the approach reduces to the traditional APCA method. One of the differences between the one-factor case in (6) and the multi-factor case in (11) is that the estimated factors and loadings in the one-factor case satisfy the conditions in (4), whereas those obtained in the multi-factor case may not. It is important to note that this discrepancy is only a finite sample outcome. Asymptotically, we can demonstrate that the multi-factor case still satisfies the identification conditions in (4). Similar outcomes are also observed in the sparse PCA literature, including Zou, Hastie, and Tibshirani (2006) and Witten, Tibshirani, and Hastie (2009), as well as in the sparse factor modeling frameworks discussed in Kristensen (2017).
For each sparse vector $\widehat{\mathbf{v}}_{i}$ obtained in Algorithm 2, the estimated factor process is obtained as $\underline{\widehat{\mathbf{f}}}_{i}=\sqrt{T}\widehat{\mathbf{v}}_{i}$. It is not hard to see that the number of iterations is $r$, and the $i$th iteration outputs a sparse estimator (multiplied by $\sqrt{T}$) of the $i$th column of ${\mathbf{F}}$. For the optimization problem in Step 4 of Algorithm 2, by the argument in Lemma A.1 of the Appendix, the matrix ${\mathbf{B}}_{i}$ in each step is a symmetric one. Therefore, we modify the truncated power method in Algorithm 1 and propose the following algorithm to obtain a sparse eigenvector for Step 4 of Algorithm 2.
Algorithm 3 (Estimation of the $i$-th eigenvector in Algorithm 2).
1:For each ${\mathbf{B}}={\mathbf{B}}_{i}$, perform a singular-value-decomposition (SVD) ${\mathbf{B}}={\mathbf{U}}{\mathbf{D}}{\mathbf{U}}^{\prime}$ and hence ${\mathbf{B}}^{1/2}={\mathbf{U}}{\mathbf{D}}^{1/2}{\mathbf{U}}^{\prime}$. Let ${\mathbf{A}}={\mathbf{B}}^{-1/2}{\mathbf{S}}{\mathbf{B}}^{-1/2}$.
2:Initialize $j=1$ and choose an initial vector ${\mathbf{x}}_{0}\in R^{N}$;
3:repeat
4:     Compute $\widetilde{\mathbf{x}}_{j}=\frac{{\mathbf{B}}^{-1/2}{\mathbf{A}}{\mathbf{x}}_{j-1}}{\|{\mathbf{A}}{\mathbf{x}}_{j-1}\|_{2}}$
5:     Let $l_{j}=\text{supp}(\widetilde{\mathbf{x}}_{j},s)$ be the indices of $\widehat{\mathbf{x}}_{t}$ with the largest $s$ absolute values.
6:     Compute ${\mathbf{x}}_{j}^{*}=\text{Truncate}(\widetilde{\mathbf{x}}_{j},l_{j})$
7:     Let ${\mathbf{x}}_{j}=\frac{{\mathbf{B}}^{1/2}{\mathbf{x}}_{j}^{*}}{\|{\mathbf{B}}^{1/2}{\mathbf{x}}_{j}^{*}\|_{2}}$.
8:     $j\leftarrow j+1$.
9:until ${\mathbf{x}}_{j}$ is convergent
10:Output: $\widehat{\mathbf{v}}_{i}={\mathbf{x}}_{j}^{*}$, which is the last ${\mathbf{x}}_{j}^{*}$ in Step 6.
Note that the symmetric matrix ${\mathbf{B}}_{i}$ in Algorithm 3 is not strictly positive definite, the half-inverse ${\mathbf{B}}_{i}^{-1/2}$ should be taken as a generalized one in line with the Moore-Penrose generalized inverse matrix.
2.6 Determination of the Number of Factors
The estimation method in Section 2.5 depends on a known number of factors $r$. In practice, $r$ is unknown and we need to develop a data-driven method to estimate it. There are several useful methods developed in the past decades to estimate the number of factors including the information criterion in Bai and Ng (2002), the random matrix theory method in Onatski (2010), the ratio-based method in Lam and Yao (2012) and Ahn and Horenstein (2013), and the white noise testing approach in Gao and Tsay (2022), among others.
In this section, we only introduce two commonly used ways, one is based on the well-known information criterion (IC) method in Bai and Ng (2002), which estimates the number of factors $r$ by
$\widehat{r}=\arg\min_{1\leq k\leq K}\log\left(\frac{1}{NT}\|{\mathbf{X}}-\widehat{\mathbf{F}}_{k}\widehat{\boldsymbol{\Lambda}}_{k}^{\prime}\|_{F}^{2}\right)+k\frac{N+T}{NT}\log(\frac{NT}{N+T}),$ (12)
where $K$ is a prescribed upper bound, $\widehat{\mathbf{F}}_{k}$ and $\widehat{\boldsymbol{\Lambda}}_{K}$ are the estimated factors and loading matrix based on $k$ factors. The other method is based on the eigenvalue ratios introduced by Lam and Yao (2012) and Ahn and Horenstein (2013). Specifically, let $\widehat{\lambda}_{1}\geq\widehat{\lambda}_{2}\geq...\geq\widehat{\lambda}_{T}$ be the $T$ sample eigenvalues of ${\mathbf{X}}{\mathbf{X}}^{\prime}$, we adopt the ratio-based method to estimate $r$ by
$\widehat{r}=\arg\min_{1\leq k\leq K}\widehat{\lambda}_{k+1}/\widehat{\lambda}_{k},$ (13)
where $K$ is a prescribed upper bound as in (12) to control the stability of the ratios. In practice, we may choose $K=\min(T,N)/3$ under the assumption that the number of factors $r$ is usually not large.
2.7 Choice of the Tuning Parameters
For a given number of factors $r$, which can be estimated by the methods in Section 2.6 above, if the cardinality of each column in ${\mathbf{F}}$ is known, we can obtain the estimated factors $\widehat{\mathbf{F}}$ by solving the optimization problem using the proposed algorithms in Section 2.5. The resulting estimated loading matrix is $\widehat{\boldsymbol{\Lambda}}^{\prime}=(\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}})^{-1}\widehat{\mathbf{F}}^{\prime}{\mathbf{X}}$, which can be obtained by the Ordinary Least-Squares (OLS) method. When the true factors are sparse and orthogonal, the theoretical results in Section 3 below suggest that $\widehat{\boldsymbol{\Lambda}}\approx{\mathbf{X}}^{\prime}\widehat{\mathbf{F}}$, implying that each column loading vector is a sparse linear combination of the panel data over the timeline.
In practice, the cardinality of each column in ${\mathbf{F}}$ is unknown, and we may choose them by the cross-validation method which is commonly used in machine learning literature. In fact, even though the sparsity parameters $s_{1},...,s_{r}$ of $r$ factor processes may not be the same, it is still a convenient way to assume that $s_{1}=...=s_{r}=s_{0}$ which can simplify the cross-validation procedure significantly. According to the theoretical results in Section 3 and the proofs in the Appendix, when the $r$ sparsity parameters are distinct, the convergence of the estimator for $s^{*}$ can be achieved, where $s^{*}=\max(s_{1},...,s_{r})$. Therefore, we assume that the sparsity parameters of each column are the same in this section. The cross-validation procedure is as follows.
For each fixed $s\in\mathbb{S}$ where $\mathbb{S}$ is a chosen candidate set for the sparsity parameter $s$, we first randomly divide the panel into two segments consisting of one training sample with $N_{1}$ components and the other a testing one with $N_{2}$ components, where $N_{1}\asymp N_{2}\asymp N$ and $N_{1}+N_{2}=N$. Equivalently, we partition the data matrix $\widetilde{\mathbf{X}}\in R^{T\times N}$ into $\widetilde{\mathbf{X}}_{1}$ and $\widetilde{\mathbf{X}}_{2}$, where $\widetilde{\mathbf{X}}_{1}\in R^{T\times N_{1}}$ and $\widetilde{\mathbf{X}}_{2}\in R^{T\times N_{2}}$. Let $\widetilde{\mathbf{F}}_{1}^{s}$ be the estimated factors based on the training sample $\widetilde{\mathbf{X}}_{1}$ and $\widetilde{\boldsymbol{\Lambda}}_{1}$ be the estimated loading matrix. Define the testing error as
$R(s,\widetilde{\mathbf{F}}_{1}^{s})=\|\widetilde{\mathbf{X}}_{2}-\widetilde{\mathbf{F}}_{1}^{s}(\widetilde{\mathbf{F}}_{1}^{s}{{}^{\prime}}\widetilde{\mathbf{F}}_{1}^{s})^{-1}\widetilde{\mathbf{F}}_{1}^{s}{{}^{\prime}}\widetilde{\mathbf{X}}_{2}\|_{F}^{2}/(N_{2}T).$ (14)
When we range the parameter $s$ over the candidate set $\mathbb{S}$, the optimal number of nonzero factors, denoted by $\widehat{s}$, is the one that produces the smallest errors in (14). On the other hand, we do not expect $s$ to be a large one to avoid the overfitting issue. Therefore, we define $g(N_{1},T)$ as a penalty function, and let
$PC(s)=R(s,\widetilde{\mathbf{F}}_{1}^{s})+rsg(N_{1},T).$
Then, $\widehat{s}$ is the one that minimizes the above criterion, that is,
$\widehat{s}=\arg\min_{s\in\mathbb{S}}PC(s)=\arg\min_{s\in\mathbb{S}}\{R(s,\widetilde{\mathbf{F}}_{1}^{s})+rsg(N_{1},T)\},$ (15)
implying that we choose the optimal $s$ which results in minimal testing errors in the cross-validations. There are many choices for the penalty function $g(N_{1},T)$ as shown in Theorem 4 in Section 3 below. We will discuss the required properties of $g(N_{1},T)$ such that we can establish the consistency of $\widehat{s}$ in theory.
In practice, we may replace the candidate set $\mathbb{S}$ by $1\leq s\leq smax$ where $smax$ is a prescribed bounded integer. Note that the cross-validation method is different from the classical one in the machine learning literature where the samples are partitioned along the time horizon, whereas we partition the samples over the cross-section in this paper. In fact, we can even perform more cross-validation experiments in order to obtain an optimal parameter $s$. For example, we can choose a sufficiently large integer $J>0$, and we can obtain $J$ testing errors as that in (14) for each fixed $s$ by partitioning the samples for $J$ times. We may denote the testing error by $R_{j}(s,\widetilde{\mathbf{F}}_{1,j}^{s})$ for the $j$-th random partition and the average testing error for $s$ is defined as
$R^{J}(s)=\frac{1}{J}\sum_{j=1}^{J}R_{j}(s,\widetilde{\mathbf{F}}_{1,j}^{s}),$ (16)
where $\widetilde{\mathbf{F}}_{1,j}^{s}$ is the estimated factors based on the training sample in the $j$-th cross-validation. Then, the optimal number of nonzero factors $\widehat{s}$ is the one that produces the smallest error of the information criterion in (15) by replacing the $R(s,\widetilde{\mathbf{F}}_{1}^{s})$ therein by $R^{J}(s)$ in Equation (16) for $s\in\mathbb{S}$.
3 Theoretical Properties
In this section, we present the asymptotic consistency and stochastic bounds of the estimators from Section 2 under under the framework where $N,T\rightarrow\infty$. Assuming the number of factors $r$ is fixed, and $s_{i}$ can be either fixed or slowly increasing, we first derive the asymptotic results under known $r$ and $s_{i}$. We also demonstrate the consistency of the estimators for $r$ and $s_{i}$ thereafter. The mathematical proofs are provided in the Appendix.
To derive the asymptotic properties of the proposed method and estimators from Section 2, we need the following high-level assumptions. These assumptions can be readily verified by some standard and more primitive assumptions, which will be discussed in detail later. Most of the assumptions below are commonly used in the PCA or approximate-factor modeling literature, and some of which are stronger than those in Bai and Ng (2002) in order to show that the estimated sparse factor $\widehat{\mathbf{F}}$ is close to ${\mathbf{F}}$. We will use $C$ or $c$ to denote a generic positive constant the value of which may change at different places.
Assumption 1.
The process $\{{\mathbf{f}}_{t}\}$ is $\alpha$-mixing with the mixing coefficients satisfying the condition $\alpha_{N}(k)<\exp(-k)$, where $\alpha_{N}(k)$ is defined as
$\alpha_{N}(k)=\sup_{i}\sup_{A\in\mathcal{F}_{-\infty}^{i},B\in\mathcal{F}_{i+k}^{\infty}}|P(A\cap B)-P(A)P(B)|,$ (17)
where $\mathcal{F}_{i}^{j}$ is the $\sigma$-field generated by $\{{\mathbf{f}}_{t}:i\leq t\leq j\}$.
Assumption 2.
Assume $E\|{\mathbf{f}}_{t}\|_{2}^{4}\leq C$ and the covariance satisfies $\frac{1}{T}\sum_{t=1}^{T}{\mathbf{f}}_{t}{\mathbf{f}}_{t}^{\prime}={\mathbf{I}}_{r}$.
Assumption 3.
There exists a vector ${\mathbf{s}}=(s_{1},...,s_{r})^{\prime}$ such that $\|{\mathbf{F}}\|_{0}\leq{\mathbf{s}}$, where ${\mathbf{F}}$ is defined as that in (3).
Assumption 4.
The eigenvalues of $\frac{1}{N}\boldsymbol{\Lambda}^{\prime}\boldsymbol{\Lambda}$ are distinct and bounded away from 00 and $\infty$ as $N\rightarrow\infty$.
Assumption 5.
$e_{i,t}=\sigma_{i}{{\varepsilon}}_{i,t}$ for some $0<c_{1}\leq\sigma_{i}\leq c_{2}<\infty$, where ${{\varepsilon}}_{i,t}$ is independent and identically distributed over $i$ and $t$.
Assumption 6.
$\{\boldsymbol{\lambda}_{i}\}$, $\{{\mathbf{f}}_{t}\}$, and $\{e_{i,t}\}$ are mutually independent groups.
Assumption 7.
$\underline{{\mathbf{f}}}_{i}$ and $\underline{{\mathbf{e}}}_{j}$ are sub-Gaussian variables for $i=1,...,r$ and $j=1,...,N$ in the sense that
$P(|{\mathbf{v}}^{\prime}\underline{{\mathbf{f}}_{i}}|>x)\leq c\exp(-cx^{2})\quad\text{and}\quad P(|{\mathbf{v}}^{\prime}\underline{{\mathbf{e}}_{i}}|>x)\leq c\exp(-cx^{2}),$
for any $\|{\mathbf{v}}\|_{2}=1$.
Assumption 1 is standard to characterize the dynamic dependence of the factor processes. See, for example, the Lemma 1 in the appendix of Gao, Ma, Wang, and Yao (2019). Assumptions 2-3 impose an identification condition and a sparsity structure on the true factor process ${\mathbf{F}}$. Assumption 4 specifies that the top eigenvalues of the covariance of the panel data are distinct in order to avoid multiplicity of eigenvalues. Assumptions 2 and 4 provide identifiability for the Model (3) so that the factors and the loading matrix can be uniquely determined asymptotically. See Bai and Ng (2013) for a detailed argument. Assumption 5 is only made to simplify the theoretical derivations. It can be relaxed to a weaker assumption such as Conditions (c.i)–(c.iii) in Bai and Ng (2013). Similar assumptions are also made in Huang, Jiang, Li, Tong, and Zhou (2022) to simplify the theoretical derivation. Assumption 7 is used to derive the Bernstein-type concentration inequality in establishing the convergence rates of the estimated factors, and it can be relaxed to much weaker ones as those in Merlevède, Peligrad, and Rio (2011).
The following theorem establishes the consistency of the first estimated factor process.
Theorem 1.
Let Assumptions 17 in the Appendix hold and $\widehat{\mathbf{v}}_{1}$ be the first solution of (6). As $N,T\rightarrow\infty$, it holds that
$\sqrt{1-(\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1})^{2}}=O_{p}\left(\sqrt{\frac{s_{1}\log(N)}{NT}}+\frac{s_{1}\log(N)}{NT}+\frac{1}{T}\right),$
where $\widehat{\mathbf{v}}_{1}=\underline{\widehat{\mathbf{f}}}_{1}/\sqrt{T}$ and ${\mathbf{v}}_{1}=\underline{{\mathbf{f}}}_{1}/\sqrt{T}$.
From Theorem 1, we see that the angle between the estimated direction $\widehat{\mathbf{v}}_{1}$ and the true one ${\mathbf{v}}_{1}$ is asymptotically equal to zero or $\pi$ if $s_{1}\log(N)/(NT)\rightarrow 0$. In fact, we can rewrite $\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1}$ as $\cos(\theta)$ where $\theta$ is the angle between $\widehat{\mathbf{v}}_{1}$ and ${\mathbf{v}}_{1}$. Some remarks for Theorem 1 are as follows
Remark 1.
(i) It follows immediately from Theorem 1 that
$\sqrt{1-\cos^{2}(\theta)}=|\sin(\theta)|=O_{p}\left(\sqrt{\frac{s_{1}\log(N)}{NT}}+\frac{s_{1}\log(N)}{NT}+\frac{1}{T}\right).$
If $s_{1}\log(N)/(NT)\rightarrow 0$, we have that $\sin(\theta)\rightarrow 0$ asymptotically, implying that $\theta\rightarrow 0$ or $\theta\rightarrow\pi$. Therefore, $\widehat{\mathbf{v}}_{1}$ and ${\mathbf{v}}_{1}$ coincide on the same line.
(ii) When the two vectors $\widehat{\mathbf{v}}_{1}$ and ${\mathbf{v}}_{1}$ point to the same direction in the sense that the angle $\theta$ is acute, then $\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1}\geq 0$ and $1\leq 1+\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1}\leq 2$. Theorem 1 implies that
$\|\widehat{\mathbf{v}}_{1}-{\mathbf{v}}_{1}\|_{2}=\sqrt{2(1-\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1})}\leq\sqrt{2(1-(\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1})^{2})}=O_{p}\left(\sqrt{\frac{s_{1}\log(N)}{NT}}+\frac{s_{1}\log(N)}{NT}+\frac{1}{T}\right),$
and therefore,
$\|\underline{\widehat{\mathbf{f}}}_{1}-\underline{{\mathbf{f}}}_{1}\|_{2}=O_{p}\left(\sqrt{\frac{s_{1}\log(N)}{N}}+\frac{s_{1}\log(N)}{N\sqrt{T}}+\frac{1}{\sqrt{T}}\right),$
implying that we can uniquely determine the factors asymptotically if $s_{1}\log(N)=o(N)$ as $N,T\rightarrow\infty$.
Next, we present the theorem concerning the consistency of all the estimated factors along the time horizon. For the distance between two matrices ${\mathbf{H}}_{1}$ and ${\mathbf{H}}_{2}$, there are several measures that can be used in the literature. For example, We adopt the discrepancy measure used by Pan and Yao (2008): for two $T\times r$ semi-orthogonal matrices ${\bf H}_{1}$ and ${\bf H}_{2}$ satisfying the condition ${\bf H}_{1}^{\prime}{\bf H}_{1}={\bf H}_{2}^{\prime}{\bf H}_{2}={\mathbf{I}}_{r}$, the difference between the two linear spaces $\mathcal{M}({\bf H}_{1})$ and $\mathcal{M}({\bf H}_{2})$ is measured by
$D(\mathcal{M}({\bf H}_{1}),\mathcal{M}({\bf H}_{2}))=\sqrt{1-\frac{1}{r}\textrm{tr}({\bf H}_{1}{\bf H}_{1}^{\prime}{\bf H}_{2}{\bf H}_{2}^{\prime})}.$ (18)
Note that $D(\mathcal{M}({\bf H}_{1}),\mathcal{M}({\bf H}_{2}))\in[0,1].$ It is equal to 00 if and only if $\mathcal{M}({\bf H}_{1})=\mathcal{M}({\bf H}_{2})$, and to $1$ if and only if $\mathcal{M}({\bf H}_{1})\perp\mathcal{M}({\bf H}_{2})$. By Lemma A1(i) in Pan and Yao (2008), $D(\cdot,\cdot)$ is a well-defined distance measure on some quotient space of matrices. Alternatively, we may also adopt the measure
$\rho({\mathbf{H}}_{1},{\mathbf{H}}_{2})=\|{\mathbf{H}}_{1}{\mathbf{H}}_{1}^{\prime}-{\mathbf{H}}_{2}{\mathbf{H}}_{2}^{\prime}\|_{F},$ (19)
which is the Frobenius norm of the difference between the projection matrices of two spaces and is also a well-defined distance between linear subspaces. In addition, if we denote the singular values of ${\mathbf{H}}_{1}^{\prime}{\mathbf{H}}_{2}$ by $\{\sigma_{i}\}_{i=1}^{r}$, in descending order, then the principal angles between $\mathcal{M}({\mathbf{H}}_{1})$ and $\mathcal{M}({\mathbf{H}}_{2})$, $\boldsymbol{\Theta}({\mathbf{H}}_{1},{\mathbf{H}}_{2})=\mbox{diag}(\theta_{1},...,\theta_{r})$, are defined as $\mbox{diag}\{\cos^{-1}(\sigma_{1}),...,\cos^{-1}(\sigma_{r})\}$; see, for example, Theorem I.5.5 of Stewart and Sun (1990). The squared Frobenius norm of the so-called $\sin\boldsymbol{\Theta}$ distance, defined as
$\|\sin\boldsymbol{\Theta}({\mathbf{H}}_{1},{\mathbf{H}}_{2})\|_{F}^{2}:=\sum_{i=1}^{r}\sin^{2}(\theta_{i}),$ (20)
can also be used to measure the distance between two linear spaces. In fact, if $r$ is finite, the distances in (18)–(20) are equivalent, because
$\rho^{2}({\mathbf{H}}_{1},{\mathbf{H}}_{2})=$ $\|{\mathbf{H}}_{1}{\mathbf{H}}_{1}^{\prime}\|_{F}^{2}+\|{\mathbf{H}}_{2}{\mathbf{H}}_{2}^{\prime}\|_{F}^{2}-2\mathrm{trace}({\mathbf{H}}_{1}{\mathbf{H}}_{1}^{\prime}{\mathbf{H}}_{2}{\mathbf{H}}_{2}^{\prime})=2r\{D(\mathcal{M}({\bf H}_{1}),\mathcal{M}({\bf H}_{2}))\}^{2}$
$=$ $2r-2\|{\mathbf{H}}_{1}^{\prime}{\mathbf{H}}_{2}\|_{F}^{2}=2\sum_{i=1}^{r}(1-\sigma_{i}^{2})=2\sum_{i=1}^{r}\sin^{2}(\theta_{i})=2\|\sin\boldsymbol{\Theta}({\mathbf{H}}_{1},{\mathbf{H}}_{2})\|_{F}^{2}.$ (21)
Therefore, we shall only use the distance in (19) to present our theoretical results in the main article.
Theorem 2.
Let Assumptions 17 in the Appendix hold and $\widehat{\mathbf{V}}$ be the matrix consisting of the estimated sparse eigenvectors obtained by Algorithm 2 . As $N,T\rightarrow\infty$, it holds that
$\rho(\widehat{\mathbf{V}},{\mathbf{V}}):=\|\widehat{\mathbf{V}}\widehat{\mathbf{V}}^{\prime}-{\mathbf{V}}{\mathbf{V}}^{\prime}\|_{F}=O_{p}\left(\sqrt{\frac{s^{*}\log(N)}{NT}}+\frac{s^{*}\log(N)}{NT}+\frac{1}{T}\right),$
where $s^{*}=\max\{s_{i}\}_{i=1}^{r}$, $\widehat{\mathbf{V}}=\widehat{\mathbf{F}}/\sqrt{T}$, and ${\mathbf{V}}={\mathbf{F}}/\sqrt{T}$.
Remark 2.
(i) From Theorem 2, we may plug the estimated factor and obtain
$\sqrt{\sum_{i=1}^{r}\sin^{2}(\theta_{i})}=O_{p}\left(\sqrt{\frac{s^{*}\log(N)}{NT}}+\frac{s^{*}\log(N)}{NT}+\frac{1}{T}\right),$
implying that all the principal angles will be asymptotically zero or $\pi$ if $s^{*}\log(N)=o(NT)$. On the other hand, if all directions of the estimated eigenvectors are in the same direction of the corresponding true ones.
(ii) Suppose the singular values of $\widehat{\mathbf{V}}^{\prime}{\mathbf{V}}$ are $\{\sigma_{i},i=1,...,r\}$ where $1\geq\sigma_{i}\geq 0$, then $\mbox{tr}(\widehat{\mathbf{V}}^{\prime}{\mathbf{V}})=\sum_{i=1}^{r}\sigma_{i}$. If all the directions of the estimated eigenvectors and the true ones coincide on the same line, by an elementary argument, we can show that
$\|\widehat{\mathbf{V}}-{\mathbf{V}}\|_{F}^{2}=\mbox{tr}[(\widehat{\mathbf{V}}-{\mathbf{V}})^{\prime}(\widehat{\mathbf{V}}-{\mathbf{V}})]=2\sum_{i=1}^{r}(1-\sigma_{i})\leq 2\sum_{i=1}^{r}(1-\sigma_{i}^{2}),$
where we used the inequality $(1-\sigma_{i})\leq(1-\sigma_{i})(1+\sigma_{i})=(1-\sigma_{i}^{2})$. Therefore, by (3), we have that
$\|\widehat{\mathbf{V}}-{\mathbf{V}}\|_{F}\asymp\rho(\widehat{\mathbf{V}},{\mathbf{V}})=O_{p}\left(\sqrt{\frac{s^{*}\log(N)}{NT}}+\frac{s^{*}\log(N)}{NT}+\frac{1}{T}\right).$
Consequently, we obtain the result for $\widehat{\mathbf{F}}$ as
$\|\widehat{\mathbf{F}}-{\mathbf{F}}\|_{F}\asymp\sqrt{T}\rho(\widehat{\mathbf{V}},{\mathbf{V}})=O_{p}\left(\sqrt{\frac{s^{*}\log(N)}{N}}+\frac{s^{*}\log(N)}{N\sqrt{T}}+\frac{1}{\sqrt{T}}\right),$
which is similar to the one in Remark 1(ii) for each single factor process.
(iii) When there is no sparsity, i.e., $s^{*}=T$, we have that
$\frac{1}{\sqrt{T}}\|\widehat{\mathbf{F}}-{\mathbf{F}}\|_{F}\asymp\rho(\widehat{\mathbf{V}},{\mathbf{V}})=O_{p}\left(\sqrt{\frac{\log(N)}{N}}+\frac{1}{T}\right),$
which is in line with the conventional result in factor modeling if we ignore the $\log(N)$ term. See Gao and Tsay (2024) for details.
Next, we establish the theoretical results for the estimated loading matrix $\widehat{\boldsymbol{\Lambda}}$ in the following theorem.
Theorem 3.
Let Assumptions 17 in the Appendix hold.
(i) There exits a rotation matrix ${\mathbf{H}}_{s}$ such that
$\max_{1\leq i\leq N}\|\widehat{\boldsymbol{\lambda}}_{i}-{\mathbf{H}}_{s}\boldsymbol{\lambda}_{i}\|_{2}=O_{p}(\sqrt{\frac{s^{*}}{NT}}\log(N)+\frac{s^{*}\log(N)^{3/2}}{NT}+\sqrt{\frac{\log(N)}{T}}),$ (22)
where ${\mathbf{H}}_{s}$ is specified in (IA.18) of the Appendix. Furthermore, if $s^{*}$ is finite, then
$\max_{1\leq i\leq N}\|\widehat{\boldsymbol{\lambda}}_{i}-{\mathbf{H}}_{s}\boldsymbol{\lambda}_{i}\|_{2}=O_{p}(\sqrt{\frac{\log(N)}{T}}).$
(ii) For $1\leq i\leq N$, if $s^{*}\log(N)=o(N)$, then there exists a rotation matrix ${\mathbf{H}}_{s}$ such that
$\sqrt{T}(\widehat{\boldsymbol{\lambda}}_{i}-{\mathbf{H}}_{s}\boldsymbol{\lambda}_{i})\longrightarrow_{d}N(0,{\mathbf{Q}}^{-1}\boldsymbol{\Gamma}_{i}{\mathbf{Q}}^{-1}),$ (23)
where ${\mathbf{Q}}$ is the limit of $\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}}/T$, and $\boldsymbol{\Gamma}_{i}=\lim_{T\rightarrow\infty}\textnormal{Var}(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}{\mathbf{f}}_{t}e_{i,t})$.
Remark 3.
By Assumptions 2 and the results in Theorem 1, it is not hard to show that ${\mathbf{Q}}={\mathbf{I}}_{r}$, but we can use the sample version $\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}}/T$ in empirical applications. If Assumption 5 holds, then the variance term in (23) reduces to $\sigma_{i}^{2}{\mathbf{I}}_{r}$, where $\sigma_{i}$ can be estimated from the residuals. Under the general setting that the noise are not $i.i.d.$, we can use the well-known heteroskedasticity and autocorrelation consistent (HAC) estimator such as that in Andrews (1991) to approximate the long-run covariance term $\boldsymbol{\Gamma}_{i}$.
Furthermore, we discuss the consistency of the estimated number of factors. Given the convergence results in Theorems 1 and 2, it is not hard to see that the estimated number of factors are consistent no matter whether we apply the information criterion of Bai and Ng (2002) or the eigenvalue ratio based method in Lam and Yao (2012) and Ahn and Horenstein (2013). The detailed proofs are the same as the techniques used therein, and we omit them to save space here.
Finally, we provide the consistency of the cross-validation method in estimating the sparsity parameters $s_{1},...,s_{r}$. We consider the case where the sparsity parameters of each column of ${\mathbf{F}}$ are the same, i.e., $s_{1}=...=s_{r}=s_{0}$, as discussed in Section 2.7. Otherwise, the following theorem states the consistency for estimating the largest sparsity parameter among the $r$ columns.
Theorem 4.
Let Assumptions 17 in the Appendix hold. Suppose $\widehat{s}$ is the solution in (15). If $g(N_{1},T)\rightarrow 0$ and $C_{N_{1}T}^{-2}g(N_{1},T)\rightarrow\infty$ as $N,T\rightarrow\infty$, we have
$\lim_{N,T\rightarrow\infty}P(\widehat{s}=s_{0})=1,$
where $s_{0}$ is the sparsity of each column of ${\mathbf{F}}$.
Remark 4.
(i) When the columns of ${\mathbf{F}}$ have distinct sparsity parameters, it is an important step to estimate the largest one among $\{s_{1},...,s_{r}\}$ first, and $\widehat{s}$ in Theorem 4 is an estimator for $s^{*}=\max\{s_{1},...,s_{r}\}$. In other words, we can show that
$\lim_{N,T\rightarrow\infty}P(\widehat{s}=s^{*})=1.$ (24)
On the other hand, estimating the largest sparsity parameter is often adequate since all the important factors over the timeline can be recovered. Furthermore, we can subsequently estimate the sparsity of each factor sequence using a coordinate descent approach.
(ii) In practice, there are many choices for $g(N_{1},T)$ in (15). For example, we may take
$g(N_{1},T)=\frac{N_{1}+T}{N_{1}T}\ln(\frac{N_{1}T}{N_{1}+T}),$
and
$IC(s)=\ln(R^{J}(s))+rs\frac{N_{1}+T}{N_{1}T}\ln(\frac{N_{1}T}{N_{1}+T})$ (25)
can consistently estimate $s_{0}$, where $r$ can be replaced by $\widehat{r}$ obtained in (13).
(iv) When $s$ is diverging with the sample size $T$ in the sense that $s/a_{T}\asymp 1$, then $s$ should be replaced by $s/a_{T}$. For example, if $s/\sqrt{T}\asymp 1$, then the information criterion in (25) will be
$IC(s)=\ln(R^{J}(s))+r\frac{s}{\sqrt{T}}\frac{N_{1}+T}{N_{1}T}\ln(\frac{N_{1}T}{N_{1}+T}),$ (26)
which works well as shown in the simulation studies.
4 Simulation Evidence
In this section, we illustrate the finite-sample properties of the proposed methodology under different choices of $N$ and $T$. To ensure our simulation results are reproducible, we set the seed to 1234 in R programming.
4.1 One-factor Case
First, we consider the one-factor case in Model (2), i.e., $r=1$ is used in this section. The factor process is generated by
$f_{t}=\phi f_{t-1}+\eta_{t},\quad\eta_{t}\overset{i.i.d.}{\sim}N(0,1),\quad t=1,...,T,$
where we set $\phi=0.5$. We consider the dimensions $N=50,100,150,300$ and $500$, with sample sizes $T=200,500,800,1000$, and $1200$ for each $N$. Each loading value is generated independently from the uniform distribution $U(-2,2)$, and the $N$-dimensional loading vector is then re-normalized to have an $l_{2}$-norm strength of $\sqrt{N}$. For each configuration of $(N,T)$, we choose the sparsity $s=\left\lceil\sqrt{T}\right\rceil$, which is the nearest integer greater than or equal to $\sqrt{T}$, and we randomly generate $s$ integers from $\{1,...,T\}$ and keep those $s$ elements in $\underline{{\mathbf{f}}}_{1}=(f_{1},...,f_{T})^{\prime}$ to be nonzero, re-normalizing the sparse vector $\underline{{\mathbf{f}}}_{1}=(f_{1},...,f_{T})^{\prime}$ to have unit variance. We consider two scenarios for the idiosyncratic terms ${\mathbf{e}}_{t}$’s in each experiment:
$(1)\,\,{\mathbf{e}}_{t}\overset{i.i.d.}{\sim}N({\bf 0},{\mathbf{I}}_{N})\quad\text{and}\quad(2)\,\,{\mathbf{e}}_{t}=\boldsymbol{\Phi}{\mathbf{e}}_{t-1}+\mbox{\boldmath$\varepsilon$}_{t},\mbox{\boldmath$\varepsilon$}_{t}\overset{i.i.d.}{\sim}N({\bf 0},{\mathbf{I}}_{N}),t=1,...,T,$
where $\boldsymbol{\Phi}$ is a diagonal matrix with the diagonal elements generated from $U(0.5,0.9)\cup U(-0.9,-0.5)$. A total of 500 replications are used throughout the experiments.
We first study the accuracy of the estimated factors by defining the estimation error in each replication as
$d(\underline{\widehat{f}}_{1},\underline{f}_{1})=\sqrt{1-(\underline{\widehat{f}}_{1}^{\prime}\underline{f}_{1}/T)^{2}}.$ (27)
[…]
Table 1: The estimation accuracy of factors when the number of factors $r=1$. The results are based on the average of the measures defined in (27) through 500 replications. $N$ and $T$ denote the dimension and the sample size, respectively.
${\mathbf{e}}_{t}\overset{i.i.d.}{\sim}N({\bf 0},{\mathbf{I}}_{N})$ ${\mathbf{e}}_{t}=\Phi{\mathbf{e}}_{t-1}+\mbox{\boldmath$\varepsilon$}_{t},\mbox{\boldmath$\varepsilon$}_{t}\overset{i.i.d.}{\sim}N({\bf 0},{\mathbf{I}}_{N})$
$T$ $T$
$N$ 200 500 800 1000 1200 200 500 800 1000 1200
50 0.050 0.040 0.036 0.034 0.032 0.087 0.069 0.054 0.051 0.055
100 0.033 0.027 0.024 0.022 0.021 0.055 0.041 0.040 0.036 0.035
150 0.026 0.021 0.018 0.017 0.016 0.044 0.035 0.031 0.030 0.028
300 0.018 0.014 0.013 0.012 0.011 0.029 0.023 0.020 0.019 0.018
500 0.014 0.011 0.009 0.009 0.008 0.022 0.017 0.015 0.014 0.013
Next, we study the accuracy of the empirical recovery (ER) of the sparse factors. Let $S$ be the set of true indexes of the $s$ nonzero elements in $\underline{f}_{1}$, and $\widehat{S}$ the set of indices of the $\widehat{s}$ nonzero elements in $\underline{\widehat{f}}_{1}$. For each replication, define the empirical recovery rate of the non-sparse indices as
$ER(s):=\frac{\#\{S\cap\widehat{S}\}}{s},$ (28)
where #$\#\{S\cap\widehat{S}\}$ is the cardinality of the intersection between the estimated indices and the true ones. Table 2 presents the empirical recovery rate of the sparsity in the factor process using the measure defined in (28). From Table 2 we see that the empirical accuracy increases as the the sample size increases for each fixed dimension $N$, and it also increases as the dimension $N$ increases for each fixed sample size $T$. This is also in agreement with our asymptotic results in Section 3.
Table 2: Empirical recovery rate of the sparsity in the factor process when the number of factors $r=1$. The empirical recovery rate in each replication is calculated as in (28). $N$ and $T$ denote the dimension and the sample size, respectively. 500 replications are used in the experiments.
${\mathbf{e}}_{t}\overset{i.i.d.}{\sim}N({\bf 0},{\mathbf{I}}_{N})$ ${\mathbf{e}}_{t}=\Phi{\mathbf{e}}_{t-1}+\mbox{\boldmath$\varepsilon$}_{t},\mbox{\boldmath$\varepsilon$}_{t}\overset{i.i.d.}{\sim}N({\bf 0},{\mathbf{I}}_{N})$
$T$ $T$
$N$ 200 500 800 1000 1200 200 500 800 1000 1200
50 0.918 0.933 0.936 0.938 0.940 0.884 0.900 0.914 0.916 0.910
100 0.940 0.950 0.954 0.956 0.957 0.916 0.930 0.930 0.935 0.937
150 0.952 0.959 0.962 0.963 0.965 0.931 0.939 0.943 0.945 0.947
300 0.969 0.971 0.971 0.974 0.973 0.951 0.956 0.958 0.962 0.961
500 0.971 0.978 0.979 0.980 0.980 0.957 0.965 0.969 0.970 0.971
Finally, we evaluate the distribution of the estimated loadings as described in Theorem 3. For simplicity, we consider the case when the noises are $i.i.d.$ and plot the empirical histogram of the estimated loading associated with the first series in Figure 1. The variance of the normal curve is based on the average of the estimated variance across 2000 replications. From Figure 1, we see that the estimators behave closely to normal, which is in agreement with our asymptotic theory.
Figure 1: Histograms of the estimated loading of $\sqrt{T}(\widehat{\boldsymbol{\lambda}}_{1}-{\mathbf{H}}_{s}\boldsymbol{\lambda}_{1})$, where we set $(N,T)=(300,500)$ and $s=\left\lceil\sqrt{T}\right\rceil$. The settings of the parameters are the same as those in Table 1. The curve represents the density of normal distribution where the standard errors are estimated by the sample versions using the residuals. 2000 replications are used in the experiments.
4.2 Multi-factor Case
In this section, we further verify the efficacy of the proposed algorithms when there are multiple factors. The number of factors is set to be $r=3$, and the factors are generated by
${\mathbf{f}}_{t}=\boldsymbol{\Psi}{\mathbf{f}}_{t-1}+\boldsymbol{\eta}_{t},\boldsymbol{\eta}_{t}\overset{i.i.d.}{\sim}N({\bf 0},{\mathbf{I}}_{r}),t=1,...,T,$
where $\Psi=\mbox{diag}(0.5,-0.6,0.7)$ is a diagonal matrix. We consider the dimensions $N=50,100,150,200$, and $300$, with the sample size $T=100,200,300,500$, and $800$ for each $N$ in this experiment. For each configuration of $(N,T)$, we set the sparsity parameters to $s_{1}=s_{2}=s_{3}=\left\lceil\sqrt{T}\right\rceil$. We first randomly generate $s_{1}$ indexes from $\{1,...,T\}$ to form an index set $S_{1}$, such that we only keep the corresponding $s_{1}$ time points of $\underline{{\mathbf{f}}}_{1}$ and the remaining ones are set to zero. The index set $S_{2}$ is formed by generating $s_{2}$ indexes $\{1,...,T\}\setminus S_{1}$, keeping only those $s_{2}$ locations of $\underline{{\mathbf{f}}}_{2}$. the sparse $\underline{{\mathbf{f}}}_{3}$ is obtained by repeating the above procedure. Then each $\underline{{\mathbf{f}}}_{j}$ is normalized to have unit variance. For the generation of the loading matrix $\boldsymbol{\Lambda}\in R^{N\times r}$, we first generate an $N\times r$ matrix ${\mathbf{M}}$ with elements independently generated from $U(-2,2)$, we then perform a singular-value decomposition on ${\mathbf{M}}$, and the left singular matrix is ${\mathbf{U}}$. The loading $\boldsymbol{\Lambda}$ is taken as ${\mathbf{U}}$ multiplied by $\sqrt{N}\mbox{diag}(3,2,1)$ on its right. We also consider the two scenarios for the idiosyncratic terms as in Section 4.1. A total of 500 replications are used throughout the experiments.
Now, we first study the estimation accuracy of the factor processes. We define the measure of the errors of the estimated factors as
$d(\widehat{\mathbf{F}},{\mathbf{F}})=\|\frac{\widehat{\mathbf{F}}\widehat{\mathbf{F}}^{\prime}}{T}-\frac{{\mathbf{F}}{\mathbf{F}}^{\prime}}{T}\|_{F}.$ (29)
The average estimation errors of the factors over 500 replications are reported in Table 3. From Table 3, we observe a similar pattern to that in Section 4.1. For each fixed dimension $N$, the estimation error generally decreases as the sample size $T$ increases. Additionally, the error decreases as the dimension $N$ increases for each fixed $T$, which is in agreement with our asymptotic theory in Theorem 4, regardless of whether the idiosyncratic terms are i.i.d. or dynamically dependent.
Table 3: The estimation accuracy of factors when the number of factors $r=3$. The measure is defined as (29). $N$ and $T$ denote the dimension and the sample size, respectively. 500 replications are used in the experiments.
${\mathbf{e}}_{t}\overset{i.i.d.}{\sim}N({\bf 0},{\mathbf{I}}_{N})$ ${\mathbf{e}}_{t}=\Phi{\mathbf{e}}_{t-1}+\mbox{\boldmath$\varepsilon$}_{t},\mbox{\boldmath$\varepsilon$}_{t}\overset{i.i.d.}{\sim}N({\bf 0},{\mathbf{I}}_{N})$
$T$ $T$
$N$ 100 200 300 500 800 100 200 300 500 800
50 0.090 0.080 0.073 0.065 0.058 0.136 0.127 0.116 0.114 0.100
100 0.059 0.054 0.049 0.043 0.039 0.099 0.090 0.075 0.068 0.063
150 0.047 0.043 0.039 0.033 0.031 0.078 0.071 0.063 0.056 0.050
200 0.041 0.036 0.032 0.029 0.026 0.068 0.061 0.055 0.049 0.043
300 0.032 0.029 0.026 0.023 0.020 0.052 0.045 0.043 0.037 0.032
Furthermore, we study the empirical recovery (ER) of the sparsity in the factor processes. Similar to the measure in (28), we define
$ER(s):=\frac{\#\{\widehat{S}_{1}\cap S_{1}\}+\#\{\widehat{S}_{2}\cap S_{2}\}+\#\{\widehat{S}_{3}\cap S_{3}\}}{3s},$ (30)
where $s=\left\lceil\sqrt{T}\right\rceil$ and $\widehat{S}_{j}$ is the the estimated nonzero locations in $\underline{\widehat{\mathbf{f}}}_{j}$ for $j=1,2$, and $3$. Table 4 reports the empirical accuracy of the sparsity locations in the factor processes when $r=3$. From Table 4, we see that the pattern is also similar to the case when $r=1$ in Section 4.1. The empirical results are in line with our asymptotic ones in the sense that the estimation accuracy will increase as the dimension or the sample size increases.
Table 4: Empirical recovery rate of the sparsity in the factor processes when the number of factors $r=3$. The error in each replication is calculated as (30). $N$ and $T$ denote the dimension and the sample size, respectively. 500 replications are used in the experiments.
${\mathbf{e}}_{t}\overset{i.i.d.}{\sim}N({\bf 0},{\mathbf{I}}_{N})$ ${\mathbf{e}}_{t}=\Phi{\mathbf{e}}_{t-1}+\mbox{\boldmath$\varepsilon$}_{t},\mbox{\boldmath$\varepsilon$}_{t}\overset{i.i.d.}{\sim}N({\bf 0},{\mathbf{I}}_{N})$
$T$ $T$
$N$ 100 200 300 500 800 100 200 300 500 800
50 0.949 0.950 0.953 0.957 0.960 0.931 0.931 0.932 0.934 0.940
100 0.966 0.965 0.965 0.969 0.971 0.947 0.948 0.952 0.955 0.958
150 0.971 0.971 0.973 0.977 0.976 0.956 0.957 0.960 0.964 0.965
200 0.973 0.974 0.977 0.977 0.980 0.959 0.959 0.964 0.966 0.970
300 0.980 0.979 0.981 0.983 0.983 0.968 0.971 0.971 0.973 0.975
Although the ratio-based method of (13) for determining the number of factors has been shown to be valid in many previous studies, we conduct an auxiliary experiment to verify its efficacy under our setting. Table 5 reports the empirical probabilities of $P(\widehat{r}=r)$ in 500 replications under the aforementioned setting. From Table 5, we see that the ratio-based method performs well, with most of the empirical probabilities being close to one. This is understandable since all the factors are strong ones. Similar results can be found in Gao and Tsay (2023).
Table 5: Empirical probabilities (EP) of determining the number of factors using the ratio-based method in (13) when the number of factors $r=3$. $N$ and $T$ denote the dimension and the sample size, respectively. 500 replications are used in the experiments.
${\mathbf{e}}_{t}\overset{i.i.d.}{\sim}N({\bf 0},{\mathbf{I}}_{N})$ ${\mathbf{e}}_{t}=\Phi{\mathbf{e}}_{t-1}+\mbox{\boldmath$\varepsilon$}_{t},\mbox{\boldmath$\varepsilon$}_{t}\overset{i.i.d.}{\sim}N({\bf 0},{\mathbf{I}}_{N})$
$T$ $T$
$N$ 100 200 300 500 800 100 200 300 500 800
50 1 1 1 1 1 1 1 1 1 1
100 1 1 1 1 1 0.998 1 1 1 1
150 1 1 1 1 1 0.998 1 1 1 1
200 1 1 1 1 1 1 1 1 1 1
300 1 1 1 1 1 1 1 1 1 1
4.3 Determining the Sparsity with Cross-Validation
In this section, we study the estimation accuracy of the information criterion discussed in Section 2.7 and Remark 4 in estimating the sparsity of the factors. For simplicity, we only consider the case when $r=1$, but similar results can be obtained for $r>1$. The generation of the factors and the data is similar to those in Section 4.1, but we only keep the largest $s=\left\lceil\sqrt{T}\right\rceil$ elements of the factor process in absolute value for each sample size $T$. We consider the dimension $N=50,100,150,200$, and $300$, and the sample size $T=100,200,300,500$, and $800$ for each $N$ in this section. For each configuration of $(N,T)$, we set the number of partitions $J=1$ for simplicity and set $N_{1}=N_{2}=N/2$ in the partition. Since $s=\sqrt{T}$ is diverging with the sample size $T$, we will use the information criterion defined in (26) to estimate the sparsity. 500 replications are used throughout the experiment. Table 6 reports the Empirical probabilities (EP) of determining the sparsity parameter using the information criterion in (26) when the number of factors $r=1$. The empirical probabilities are calculated based on the 500 experiments. From Table 6, we see that the proposed information criterion works sufficiently well in estimating the sparsity parameters. When the idiosyncratic terms possess some dynamic dependence structure, the proposed method does not provide a probability close to one to estimate the sparsity, but the performance of the proposed method still improves as the sample size $T$ and/or the dimension $N$ increases, which is in agreement with our asymptotic theory in Theorem 4.
Table 6: Empirical probabilities (EP) of determining the sparsity parameter using the information criterion in (26) when the number of factors $r=1$. $N$ and $T$ denote the dimension and the sample size, respectively. 500 replications are used in the experiments.
${\mathbf{e}}_{t}\overset{i.i.d.}{\sim}N({\bf 0},{\mathbf{I}}_{N})$ ${\mathbf{e}}_{t}=\Phi{\mathbf{e}}_{t-1}+\mbox{\boldmath$\varepsilon$}_{t},\mbox{\boldmath$\varepsilon$}_{t}\overset{i.i.d.}{\sim}N({\bf 0},{\mathbf{I}}_{N})$
$T$ $T$
$N$ 100 200 300 500 800 100 200 300 500 800
50 1 1 1 1 1 0.882 0.980 1 1 1
100 1 1 1 1 1 0.998 1 1 1 1
150 1 1 1 1 1 1 1 1 1 1
200 1 1 1 1 1 1 1 1 1 1
300 1 1 1 1 1 1 1 1 1 1
5 An Empirical Application to Stock Returns
5.1 Data
In this section, we estimate the sparse latent risk factors across the time horizons in daily returns of individual stocks. The nonzero factors over certain time period can provide us one way to bridge time and certain type of risks in the financial market. The daily returns are downloaded from the CRSP daily security database and adjusted for dividend and stock splits. The data set used is the same as that in Pelger (2020) and consists of the daily stock returns for the balanced panel of S&P 500 stocks from January 1st 2004 to December 31st 2016. The daily interest rates from Prof. Kenneth French’s website (https://mba.tuck.dartmouth.edu/pages/faculty/ken.french/data_library.html) are used to adjust the daily returns of individual stocks. The full data set is also available at https://mpelger.people.stanford.edu/data-and-code, where only the stocks with returns available for the full-time horizon are included, leaving us with a panel of $N=332$ and $T=3273$.
Similar to Pelger (2020), we group these 332 stocks into 14 categories by industry sector, as shown in Table 7. The data encompasses a wide range of industry sectors within the stock market, including Oil, Finance, Electricity, Technology, Food, Manufacturing, Pharma&Chemicals, Primary Manufacturing, Machinery, Health, Transportation, Trade, Services, and Mining.
Table 7: Categorization of 332 Stocks in the S&P 500 Index by Industry Sector. The data spans from January 1, 2004, to December 31, 2016.
Category Stock Ticker
Oil APA, APC, CHK, CNX, COP, CVX, HAL, HP, MRO, MUR, NBR, NE, NFG, NFX, OXY, PXD, RDC, RIG, SLB, SRE, TSO, VLO, XOM
Finance ACAS, AET, AFL, AIG, AIV, AJG, ALL, AMG, AMT, AXP, BAC, BEN, BK, BXP, C, CI, CINF, CME, COF, FII, FITB, GS, HBAN, HCN, HCP, HIG, HRB, JNS, JPM, KEY, LNC, MBI, MET, MMC, MTG, NTRS, PFG, PLD, PNC, PRU, RF, RJF, SLM, SNV, SPG, SPY, STI, STT, TMK, TROW, UDR, UNM, USB, WFC, XL, ZION
Electricity AEP, AES, CMS, CNP, FE, LNT, NI, PCG, PEG, PNW, PPL, SO, XEL
Technology AAPL, ADBE, ADI, ADP, ADSK, AKAM, AMD, AMZN, APH, CA, CERN, CSC, CSCO, CTSH, CTXS, CVG, FFIV, FISV, HAR, HPQ, IBM, INTC, INTU, JBL, JNPR, KLAC, LLL, LLTC, LMT, MCHP, MSFT, MU, NCR, NTAP, NVDA, ORCL, QCOM, SNPS, STX, SWKS, SYMC, TXN, UIS, VRSN, WDC, XLNX, XRX, YHOO
Food CAG, CCE, CPB, GIS, HSY, K, KO, MKC, MO, PEP
Manufacturing CTAS, GCI, IP, JCI, KMB, LEG, LPX, MAS, NYT, RL, WY
Pharma & Chemicals ABT, AMGN, APD, AVP, BMY, CL, CLX, ENDP, GILD, IFF, JNJ, LLY, MON, MRK, MYL, OLN, PFE, PG, PPG, PRGO, PX
Primary manufacturing AA, AKS, ATI, COH, CTB, GLW, GT, NKE, NWL, OI, SNA, SWK, TUP, USG, UTX, WOR
Machinery A, AMAT, AME, BA, BC, BGG, BHI, CAT, CMI, COL, FLIR, FOSL, FTI, GD, GE, GRMN, HAS, HON, HRS, IR, ITT, LRCX, MAT, NOC, PBI, PCAR, PKI, ROK, RTN, TEN, TER, TKR, TMO, TXT, WHR
Health BAX, BCR, BDX, MDT, MMM, SYK, VAR, XRAY
Transportation ALK, CCI, CCL, CMCSA, CSX, CTL, DISH, EXPD, JBHT, KSU, LUV, LVLT, NSC, TDW, UNP, UPS, VIAB, VZ
Trade ABC, ANF, AZO, BBBY, BBY, CAH, COST, CVS, DLTR, FAST, FL, GPC, GPS, GWW, HD, HSIC, JCP, JWN, KMX, KR, KSS, LOW, MCD, MCK, ORLY, PDCO, RAD, ROST, SBUX, SHW, SPLS, SVU, SYY, TGT, TIF, TSCO, URBN, WEN, WMT, YUM
Services ACN, ADS, APOL, EBAY, ESRX, FLR, IGT, INCY, IPG, IT, KBH, LEN, LH, MAR, MCO, NFLX, OMC, PAYX, PCLN, PHM, PWR, R, RCL, RHI, SEE, SRCL, THC, TWX, UHS, URI, WYNN
Mining NEM, VMC
5.2 Sparse Factor Estimation
We first determine the number of factors for the centered data ${\mathbf{X}}$ using the eigenvalue-ratio method in (13). Figure 2 plots the ratios of eigenvalues of ${\mathbf{X}}{\mathbf{X}}^{\prime}$ and shows that the largest gap between the eigenvalues occurs between $\widehat{\lambda}_{1}$ and $\widehat{\lambda}_{2}$, implying that the number of factors is $\widehat{r}=1$. Next, we apply Algorithm 1 and the cross-validation method in Section 2.7 with the number of random partitions $J=10$, where the seed number is set to be 1234 in R. For the sparsity of the factors over the time horizon, we conduct a grid search over the time interval $s\in[\lceil\sqrt{T}\rceil-10,\lceil\sqrt{T}\rceil+150]=[48,208]$, where $\lceil x\rceil$ denotes the smallest integer that exceeds $x$. The estimated sparsity parameter is $\widehat{s}=184$, indicating that there are 184 days during which the stocks were driven by significant systematic risks. We can see that the method achieves a substantial dimension-reduction compared to the full time length $T=3273$.
Figure 2: (a) The first $10$ eigenvalues of ${\mathbf{X}}{\mathbf{X}}^{\prime}$; (b) The plots of the ratios for the eigenvalues $\widehat{\lambda}_{i}$ of the matrix $\widehat{\mathbf{S}}$.
We then plot the estimated sparse factor returns in Figure 4. The plot reveals roughly three clusters of time periods associated with significant risks that systematically affect the stock market. The first cluster is centered around 2008, driven by the 2007–2008 financial crisis. The second and third clusters are closely related, both associated with the European sovereign debt crisis. The most prominent cluster is centered around 2008, driven by the 2007–2008 financial crisis. This period saw severe disruptions in the financial markets, with significant declines in stock returns across the board. For the two closely related clusters associated with the European sovereign debt crisis, the first cluster corresponds to the initial phase of the crisis in 2010, while the second cluster relates to the intensification of the crisis in 2011 and 2012.
Next, we estimate the loadings of the 332 stock returns given the sparse common factors. We present the estimated loadings, indicating the level of dependence of 14 different industry sectors on the common risk factor in Figure 3. The loadings quantify the extent to which each sector is influenced by these common factors. Here are the detailed observations based on the average dependence values: Oil Sector: Exhibits moderate dependence on common risk factors, with an average loading of 0.0122. Finance Sector: Shows significant dependence, with an average loading of 0.0163. Electricity Sector: Demonstrates relatively low dependence, with an average loading of 0.0061. Technology Sector: Reflects moderate dependence, with an average loading of 0.0094. Food Sector: Indicates low dependence on common factors, with an average loading of 0.0044. Manufacturing Sector: Displays moderate dependence, with an average loading of 0.0113. Pharma & Chemicals Sector: Shows relatively low dependence compared to other sectors, with an average loading of 0.0068. Primary Manufacturing Sector: Exhibits moderate dependence, with an average loading of 0.0130. Machinery Sector: Demonstrates moderate dependence, with an average loading of 0.0106. Health Sector: Indicates relatively low dependence, with an average loading of 0.0059. Transportation Sector: Shows moderate dependence, with an average loading of 0.0096. Trade Sector: Reflects moderate dependence, with an average loading of 0.0081. Services Sector: Exhibits moderate dependence, with an average loading of 0.0108. Mining Sector: Indicates relatively low dependence on common risk factors, with an average loading of 0.0077.
Overall, the Finance sector shows the highest average dependence on common risk factors, while sectors like Food, Health, and Electricity exhibit lower average dependence. These variations highlight the differing levels of sensitivity across industry sectors to common economic and market risk factors.
Figure 3: Barplots of the the estimated loadings, indicating the level of dependence of 14 different industry sectors on common risk factors.
5.3 Bridging Time and Risk Factors
To further identify the systematic factors over time, referred to as time factors, Table LABEL:Table-ft in the Appendix provides a comprehensive list of dates with significant systematic risk factors, the reasons for these factors, and their associated time factors. Each entry explains the specific event and its impact on stock returns, summarized according to the descriptions of the time factors in Table 8. These reasons are extracted from the daily reports on CNN Money (www.money.cnn.com) after the market closes on each trading day.
From Table LABEL:Table-ft, we see that specific events impact the stock market on each date with identified systematic factors. We further analyze the words generated from the columns of "Reason" in Table LABEL:Table-ft through textual analysis. The word cloud shown in Figure 5 provides a visual representation of the most frequent terms and phrases mentioned as influencing factors. These words are associated with nine significant factors outlined in Table 8 that affect stock returns. We treat Credit risk, Oil price, China, and Europe as individual factors since they significantly influence stock returns during the period from January 1, 2004, to December 31, 2016. The nine factors are summarized in Table 8.
Figure 4: Time plots of the estimated sparse risk factors across the time horizon. The data consists of the daily stock returns for the balanced panel of S&P 500 stocks from January 1st 2004 to December 31st 2016.
Table 8: The nine significant systematic risks over the the time horizon from January 1 2004 to December 31, 2016, identified by the proposed sparse factor modeling approach.
Time Factor Description
Economic Indicators Economic activities such as GDP growth, inflation, interest rates, and consumer confidence, sector reports, and job reports, etc.
Government Policies Government policies, such as interest rate policies, tax policies, trade policies, and monetary policies, etc.
Global Events Global events such as political instability, terrorist attacks, natural disasters, and pandemics, etc.
Market Sentiment Market sentiment, which refers to the overall mood of investors.
Company-specific Factors Company-specific factors, such as financial performance, management quality, and competitive position, can also impact financial markets.
Credit Risk Credit risk refers to the possibility of borrowers defaulting on their obligations, affecting lenders and investors.
Oil Price A reference price for buyers and sellers of crude oil such as West Texas Intermediate (WTI), Brent Crude, Dubai Crude, OPEC Reference Basket, Tapis crude, Bonny Light, Urals oil, Isthmus, and Western Canadian Select (WCS).
China All significant economic activities associated with China such as a big decline in its financial market or a slowdown in its economic growth.
Europe All significant economic activities associated with Europe and major European countries such as a debt crisis and the Brexit.
We also plot the frequency charts of the nine factors over the time horizon in Figure 6. The most frequently mentioned factor is market sentiment, highlighting its dominant role in driving stock price fluctuations. Economic indicators and government policies follow closely, indicating the significant impact of macroeconomic data and policy decisions on market behavior. Company-specific factors also play an important role, affecting individual stock performance and, by extension, broader market trends. Factors related to Europe, such as economic conditions and crises, have a moderate influence, reflecting the interconnectedness of global markets. Oil price fluctuations, China’s economic activities, and global events are also notable contributors, underscoring the importance of global economic dynamics. Credit risk, while mentioned less frequently, remains a critical factor during periods of financial instability.
Figure 5: Word cloud generated from the column of Reason in Table LABEL:Table-ft. The data consists of the daily stock returns for the balanced panel of S&P 500 stocks from January 1st 2004 to December 31st 2016.
Figure 6: Frequency of Time Factors Influencing Stock Market Movements from January 2004 to December 2016. The chart visualizes the occurrence of various time factors, including market sentiment, economic indicators, government policies, company-specific factors, Europe, oil price, China, global events, and credit risk, based on their frequency of mention in Table LABEL:Table-ft.
5.4 Implications
From Tables LABEL:Table-ft and 8, we can see that these factors are all important to the stability of the financial market. In general, they affect the financial market in the following ways:
  1. (1)
    Economic Indicators: This includes employment reports, GDP growth, inflation rates, interest rates, and consumer confidence indices. For example, the weaker-than-expected December jobs report on January 4, 2008, exacerbated worries about a potential recession, leading to significant negative returns.
  2. (2)
    Government Policies: These involve fiscal and monetary policy decisions such as interest rate cuts, bailouts, and tax policies. For instance, the Federal Reserve’s emergency interest rate cut on January 23, 2008, initially had a mixed impact but ultimately contributed to a market bounce back.
  3. (3)
    Market Sentiment: Refers to the overall mood and confidence of investors. Market sentiment was notably affected on October 28, 2008, when investors dove back into stocks despite the ongoing financial turmoil, reflecting a temporary boost in confidence.
  4. (4)
    Company-specific Factors: Events specific to individual companies, such as financial performance, management changes, and competitive positioning, can significantly impact their stock prices and, by extension, the broader market.
  5. (5)
    Credit Risk: This pertains to the risk of defaults and financial instability within the banking sector. Renewed credit market fears on November 7, 2007, following concerns about Bear Stearns, led to a significant drop in stock returns.
  6. (6)
    Oil Price: Fluctuations in crude oil prices directly impact stock returns. For example, a spike in oil prices on June 6, 2008, contributed to significant negative returns, highlighting the sensitivity of the market to energy costs.
  7. (7)
    Global Events: Includes political instability, terrorist attacks, natural disasters, and pandemics. The news of explosions at the Boston Marathon on April 15, 2013, resulted in a sharp market decline.
  8. (8)
    China: Economic activities and market movements within China. The big decline in Chinese stocks on February 27, 2007 had a notable impact on global markets.
  9. (9)
    Europe: Economic activities and crises within Europe, including the debt crisis and Brexit. The downgrade of Greece’s debt rating to junk status on April 27, 2010, sparked fears about the European debt crisis, leading to significant negative returns.
From Tables LABEL:Table-ft and 8, we have the following findings and implications:
  1. (1)
    Understanding Systematic Risks: Identifying the key factors that drive systematic risks provides valuable insights for investors and policymakers. By recognizing these factors, they can make more informed decisions to mitigate risks and capitalize on potential opportunities.
  2. (2)
    Temporal Patterns and Market Sensitivity: The clustering around the 2008 financial crisis and the European debt crisis underscores how prolonged economic uncertainties impact market stability. The varying reactions to government policies and economic indicators suggest the critical role of investor confidence during crises. For example, the mixed market reaction to the Federal Reserve’s interest rate cut on January 23, 2008, illustrates the complexity of market sentiment during turbulent periods.
  3. (3)
    Policy Implications: Our findings emphasize the importance of timely and effective government interventions, such as interest rate cuts and bailouts, in stabilizing the markets. Clear communication from policymakers can significantly influence investor sentiment and market outcomes. The market’s positive response to the Federal Reserve’s emergency interest rate cut on January 23, 2008, and the subsequent bounce back in stock prices highlight the effectiveness of such interventions.
  4. (4)
    Sector-Specific Insights: The financial sector’s vulnerability to systemic risks, particularly credit market fears and bank-specific news, is evident. Additionally, fluctuations in oil prices impact not only the energy sector but also broader market sentiments. The significant impact of renewed credit market fears on November 7, 2007, following concerns about Bear Stearns, underscores the interconnectedness of the financial sector and overall market stability.
  5. (5)
    Global Interconnectedness: The influence of events in China and Europe on US stock returns highlights the interconnected nature of global markets. This interconnectedness suggests that investors may benefit from diversifying their portfolios to mitigate risks associated with specific regions or events. The market’s reaction to the downgrade of Greece’s debt rating on April 27, 2010, and the subsequent fears about the European debt crisis illustrate the global ripple effects of regional economic events.
Our empirical analysis using sparse factor modeling provides valuable insights into the significant risk factors influencing stock returns over time. The identified clusters of risk periods and the associated factors offer a deeper understanding of market dynamics and the impact of various economic and political events. The findings underscore the importance of timely government interventions, clear communication from policymakers, and the need for diversification in investment portfolios to manage global risks effectively.
6 Conclusion
Sparse factor modeling focuses on extracting meaningful insights by identifying a subset of relevant variables, thereby enhancing interpretability and reducing computational demands. This study proposes a novel perspective by assuming that factor processes are sparse over the timeline, rather than enforcing sparsity on loadings. This approach is particularly relevant for scenarios where systematic co-movements are significant only during certain periods, such as financial crises or policy changes.
The paper introduces approximate factor models and their estimation using conventional PCA, followed by the proposed sparse asymptotic PCA framework. It details the estimation procedures and theoretical properties. Simulation and empirical studies demonstrate the method’s effectiveness in providing interpretable results and linking co-movements to specific events. This study not only advances the methodology of sparse factor modeling but also bridges the gap between time-specific events and risk factors in economic and financial systems. The proposed approach offers a robust framework for future research, including the expansion of time horizons, sector-specific analyses, and the integration of advanced modeling techniques.
A Mathematical Proofs
We will use $C$ or $c$ to denote a generic constant the value of which may change at different places. We first introduce several notations used in the derivations. Let ${\mathbf{V}}={\mathbf{F}}/\sqrt{T}$, and ${\mathbf{v}}_{i}=\underline{{\mathbf{f}}}_{i}/\sqrt{T}$ be the $i$-th columns of ${\mathbf{V}}$. Define ${\mathbf{S}}_{c}={{\mathbf{F}}\boldsymbol{\Lambda}^{\prime}\boldsymbol{\Lambda}^{\prime}{\mathbf{F}}}/{(NT)}$ and ${\mathbf{S}}_{e}={\mathbf{e}}{\mathbf{e}}^{\prime}/(NT)$. It follows from Model (3) and Assumptions 24 that
${\mathbf{S}}={\mathbf{S}}_{c}+{\mathbf{S}}_{R}=\lambda_{1}{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}+...+\lambda_{r}{\mathbf{v}}_{r}{\mathbf{v}}_{r}^{\prime}+{\mathbf{S}}_{e}+\boldsymbol{\Sigma}_{fe}+\boldsymbol{\Sigma}_{fe}^{\prime},$ (IA.1)
where ${\mathbf{S}}_{R}={\mathbf{S}}_{e}+\boldsymbol{\Sigma}_{fe}+\boldsymbol{\Sigma}_{fe}^{\prime}$ and $\boldsymbol{\Sigma}_{fe}={\mathbf{F}}\boldsymbol{\Lambda}^{\prime}{\mathbf{e}}^{\prime}/(NT)$ We need a few lemmas first.
Lemma A.1.
For the $i$-th iteration in Algorithm 2 of the main article, the matrix ${\mathbf{B}}_{i}$ in step 7 of Algorithm 2 is symmetric for each $i=1,...,r+1$.
Proof. We will prove the result by an induction method. For each iteration $i=1,...,r+1$, we will first show that
${\mathbf{B}}_{i}^{2}={\mathbf{B}}_{i}\quad\text{and}\quad{\mathbf{B}}_{i}^{\prime}{\mathbf{B}}_{i}={\mathbf{B}}_{i},$ (IA.2)
which is stronger than the result in Lemma A.1 and implies that ${\mathbf{B}}_{i}$ is symmetric.
Note that ${\mathbf{B}}_{0}={\mathbf{I}}_{T}$ which obviously satisfies the identities in (IA.2). From Algorithm 2, we see that ${\mathbf{B}}_{1}={\mathbf{I}}_{T}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}$ which also satisfies (IA.2) since $\widehat{\mathbf{v}}_{1}^{\prime}\widehat{\mathbf{v}}_{1}=1$. Next, we assume that ${\mathbf{B}}_{j}^{2}={\mathbf{B}}_{j}$ and ${\mathbf{B}}_{j}^{\prime}{\mathbf{B}}_{j}={\mathbf{B}}_{j}$, then, by the fact that $\widehat{\mathbf{v}}_{j}^{\prime}{\mathbf{B}}_{j}\widehat{\mathbf{v}}_{j}=1$ and an elementary argument, we can show that
${\mathbf{B}}_{j+1}^{2}={\mathbf{B}}_{j}^{2}-2{\mathbf{B}}_{j}{\mathbf{q}}_{j}{\mathbf{q}}_{j}^{\prime}+{\mathbf{B}}_{j}{\mathbf{q}}_{j}\widehat{\mathbf{v}}_{j}^{\prime}{\mathbf{B}}_{j}{\mathbf{v}}_{j}{\mathbf{v}}_{j}^{\prime}{\mathbf{B}}_{j}^{\prime}={\mathbf{B}}_{j}({\mathbf{I}}_{T}-{\mathbf{q}}_{j}{\mathbf{q}}_{j}^{\prime})={\mathbf{B}}_{j+1},$
and
${\mathbf{B}}_{j+1}^{\prime}{\mathbf{B}}_{j+1}={\mathbf{B}}_{j}-2{\mathbf{B}}_{j}{\mathbf{q}}_{j}{\mathbf{q}}_{j}^{\prime}+{\mathbf{B}}_{j}{\mathbf{q}}_{j}\widehat{\mathbf{v}}_{j}^{\prime}{\mathbf{B}}_{j}\widehat{\mathbf{v}}_{j}{\mathbf{q}}_{j}^{\prime}={\mathbf{B}}_{j+1},$
where we used the assumption in (IA.2) for $i=j$ and the fact that $\widehat{\mathbf{v}}_{j}^{\prime}{\mathbf{B}}_{j}\widehat{\mathbf{v}}_{j}=1$. Therefore, this proves (IA.2) for all $i$. It follows from (IA.2) that ${\mathbf{B}}_{i}$ is a symmetric matrix for all $i=1,...,r+1$. This completes the proof. $\Box$
Lemma A.2.
For any two unit vectors ${\mathbf{x}}$, ${\mathbf{y}}\in R^{N}$, the singular values of ${\mathbf{x}}{\mathbf{x}}^{\prime}-{\mathbf{y}}{\mathbf{y}}^{\prime}$ are
$\sqrt{1-({\mathbf{x}}^{\prime}{\mathbf{y}})^{2}},\sqrt{1-({\mathbf{x}}^{\prime}{\mathbf{y}})^{2}},0,...,0,$
where there are $N-2$ zeros.
Proof. It is a 1-dimensional case of Proposition 2.1 in Vu and Lei (2013). See also the descriptions for (3.3)–(3.4) in Gao and Tsay (2021). Specifically, note that the only non-zero singular value of ${\mathbf{x}}{\mathbf{x}}^{\prime}({\mathbf{I}}_{N}-{\mathbf{y}}{\mathbf{y}}^{\prime})$ is
$\sqrt{\mbox{tr}({\mathbf{x}}{\mathbf{x}}^{\prime}({\mathbf{I}}_{N}-{\mathbf{y}}{\mathbf{y}}^{\prime})^{2}{\mathbf{x}}{\mathbf{x}}^{\prime})}=\sqrt{\mbox{tr}\{1-({\mathbf{x}}^{\prime}{\mathbf{y}})^{2}\}}=1-({\mathbf{x}}^{\prime}{\mathbf{y}})^{2}.$
By Proposition 2.1 of Vu and Lei (2013), the singular values of ${\mathbf{x}}{\mathbf{x}}^{\prime}-{\mathbf{y}}{\mathbf{y}}^{\prime}$ are
$\sqrt{1-({\mathbf{x}}^{\prime}{\mathbf{y}})^{2}},\sqrt{1-({\mathbf{x}}^{\prime}{\mathbf{y}})^{2}},0,...,0,$
where there are $N-2$ zeros. This completes the proof. $\Box$
Lemma A.3.
For any two matrices ${\mathbf{M}},{\mathbf{N}}\in R^{N}$, we have the following inequality,
$\mbox{tr}({\mathbf{M}}^{\prime}{\mathbf{N}})\leq\|{\mathbf{M}}\|_{2}\|{\mathbf{N}}\|_{*},$
where $\|{\mathbf{N}}\|_{*}$ denotes the sum of the singular values of ${\mathbf{N}}$.
Proof. We perform a singular value decomposition to ${\mathbf{N}}$ and obtain ${\mathbf{N}}={\mathbf{U}}{\mathbf{D}}{\mathbf{V}}^{\prime}$. Then,
$\mbox{tr}({\mathbf{M}}^{\prime}{\mathbf{N}})=\mbox{tr}({\mathbf{V}}^{\prime}{\mathbf{M}}^{\prime}{\mathbf{U}}{\mathbf{D}})=\sum_{i=1}^{N}({\mathbf{V}}^{\prime}{\mathbf{M}}^{\prime}{\mathbf{U}})_{ii}{\mathbf{D}}_{ii}\leq$ $\max_{i}|({\mathbf{V}}^{\prime}{\mathbf{M}}^{\prime}{\mathbf{U}})_{ii}|\sum_{i=1}^{N}{\mathbf{D}}_{ii}$
$\leq$ $\|{\mathbf{V}}^{\prime}{\mathbf{M}}^{\prime}{\mathbf{U}}\|_{2}\|{\mathbf{D}}\|_{*}$
$\leq$ $\|{\mathbf{M}}\|_{2}\|{\mathbf{N}}\|_{*}.$ (IA.3)
This completes the proof. $\Box$
Lemma A.4.
Under Assumptions 17, the estimator $\widehat{\mathbf{v}}_{1}$ of ${\mathbf{v}}_{1}$ satisfies
$\sqrt{1-(\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1})^{2}}\leq\frac{2}{\lambda_{1}-\lambda_{2}}\sup_{{\mathbf{v}}\in S_{1}}|{\mathbf{v}}^{\prime}{\mathbf{S}}_{R}{\mathbf{v}}|,$
where ${\mathbf{S}}_{R}$ is defined as that in (IA.1) and $S_{1}=\{{\mathbf{v}}:\|{\mathbf{v}}\|_{2}=1\,\,\text{and}\,\,\|{\mathbf{v}}\|_{0}\leq 2s_{1}\}$.
Proof. We only focus on the case that $\widehat{\mathbf{v}}_{1}\neq{\mathbf{v}}_{1}$ as Lemmas A.4 obviously holds in such a case. For any two vectors ${\mathbf{x}}$ and ${\mathbf{y}}$, we know the trace has the property that $\mbox{tr}({\mathbf{x}}{\mathbf{y}}^{\prime})=\mbox{tr}({\mathbf{y}}^{\prime}{\mathbf{x}})$. Note that ${\mathbf{S}}_{c}{\mathbf{v}}_{1}=\lambda_{1}{\mathbf{v}}_{1}$, then
${\mathbf{S}}_{c}-\lambda_{1}{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}=({\mathbf{I}}_{n}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}){\mathbf{S}}_{c}({\mathbf{I}}_{N}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}).$ (IA.4)
On the other hand, for any $\|{\mathbf{v}}\|_{2}=1$,
$\mbox{tr}({\mathbf{S}}_{c}({\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}-{\mathbf{v}}{\mathbf{v}}^{\prime}))=$ $\mbox{tr}({\mathbf{v}}_{1}^{\prime}{\mathbf{S}}_{c}{\mathbf{v}}_{1})-\mbox{tr}({\mathbf{v}}^{\prime}({\mathbf{S}}_{c}-\lambda_{1}{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}){\mathbf{v}})-\lambda_{1}\mbox{tr}({\mathbf{v}}^{\prime}{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}{\mathbf{v}})$
$=$ $\lambda_{1}-\lambda_{1}({\mathbf{v}}^{\prime}{\mathbf{v}}_{1})^{2}-{\mathbf{v}}^{\prime}({\mathbf{I}}_{N}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}){\mathbf{S}}_{c}({\mathbf{I}}_{N}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}){\mathbf{v}},$ (IA.5)
where we use Equation (IA.4) in the last step.
Let ${\mathbf{b}}=({\mathbf{I}}_{N}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}){\mathbf{v}}/\|({\mathbf{I}}_{N}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}){\mathbf{v}}\|_{2}$, it follows that ${\mathbf{b}}^{\prime}{\mathbf{v}}_{1}=0$ and therefore, ${\mathbf{b}}\in\text{span}\{{\mathbf{v}}_{2},...,{\mathbf{v}}_{N}\}$. By (IA.1), it follows that ${\mathbf{b}}^{\prime}{\mathbf{S}}_{c}{\mathbf{b}}\leq\lambda_{2}$, which implies that
${\mathbf{v}}^{\prime}({\mathbf{I}}_{N}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}){\mathbf{S}}_{c}({\mathbf{I}}_{N}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}){\mathbf{v}}\leq\lambda_{2}\|({\mathbf{I}}_{n}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}){\mathbf{v}}\|_{2}^{2}=\lambda_{2}-\lambda_{2}({\mathbf{v}}^{\prime}{\mathbf{v}}_{1})^{2}.$ (IA.6)
By (A) and (IA.6), we replace ${\mathbf{v}}$ by the estimator $\widehat{\mathbf{v}}_{1}$ of ${\mathbf{v}}_{1}$ and obtain
$\mbox{tr}({\mathbf{S}}_{c}({\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}))\geq(\lambda_{1}-\lambda_{2})\{1-(\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1})^{2}\}.$ (IA.7)
Now turn to the original optimization problem. Note that $\|\widehat{\mathbf{v}}_{1}\|_{2}\leq s_{1}$ and $\|{\mathbf{v}}_{1}\|_{2}\leq s_{1}$, and $\widehat{\mathbf{v}}_{1}$ is the one that maximizes the variance ${\mathbf{v}}^{\prime}{\mathbf{S}}{\mathbf{v}}$ with the constraint $\|{\mathbf{v}}\|_{2}\leq s_{1}$, then
$\mbox{tr}({\mathbf{S}}(\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}))=\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{S}}\widehat{\mathbf{v}}_{1}-{\mathbf{v}}_{1}^{\prime}{\mathbf{S}}{\mathbf{v}}_{1}>0$ (IA.8)
Combining (IA.7) and (IA.8) gives
$1-(\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1})^{2}\leq\frac{1}{\lambda_{1}-\lambda_{2}}\mbox{tr}(({\mathbf{S}}-{\mathbf{S}}_{c})(\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}))=\frac{1}{\lambda_{1}-\lambda_{2}}\mbox{tr}({\mathbf{S}}_{R}(\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime})).$ (IA.9)
Note that there are only $s_{1}$ nonzero elements in $\widehat{\mathbf{v}}_{1}$ and ${\mathbf{v}}_{1}$, and the indexes of their non-zero elements are not necessarily identical, we define a diagonal matrix ${\mathbf{D}}_{1}$ where $i$-th diagonal entry is $1$ when the corresponding $i$th element in $\widehat{\mathbf{v}}_{1}$ or ${\mathbf{v}}_{1}$ are nonzero. Then ${\mathbf{D}}_{1}\widehat{\mathbf{v}}_{1}=\widehat{\mathbf{v}}_{1}$ and ${\mathbf{D}}_{1}{\mathbf{v}}_{1}={\mathbf{v}}_{1}$, and
$\mbox{tr}(({\mathbf{S}}_{R}(\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime})=\mbox{tr}({\mathbf{S}}_{R}{\mathbf{D}}_{1}(\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}){\mathbf{D}}_{1})=\mbox{tr}({\mathbf{D}}_{1}{\mathbf{S}}_{R}{\mathbf{D}}_{1}(\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime})).$
By Lemma A.3 and Lemma A.2,
$\mbox{tr}({\mathbf{D}}_{1}{\mathbf{S}}_{R}{\mathbf{D}}_{1}(\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}))\leq\|{\mathbf{D}}_{1}{\mathbf{S}}_{R}{\mathbf{D}}_{1}\|_{2}\|\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}\|_{*}=\|{\mathbf{D}}_{1}{\mathbf{S}}_{R}{\mathbf{D}}_{1}\|_{2}2\sqrt{1-(\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1})^{2}}.$
It follows from (IA.9) and the above one that
$\sqrt{1-(\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1})^{2}}\leq\frac{2}{\lambda_{1}-\lambda_{2}}\|{\mathbf{D}}_{1}{\mathbf{S}}_{R}{\mathbf{D}}_{1}\|_{2}\leq\frac{2}{\lambda_{1}-\lambda_{2}}\sup_{\|{\mathbf{v}}\|_{2}=1}|{\mathbf{v}}^{\prime}{\mathbf{D}}_{1}{\mathbf{S}}_{R}{\mathbf{D}}_{1}{\mathbf{v}}|.$ (IA.10)
Note that ${\mathbf{D}}_{1}{\mathbf{v}}$ have at most $2s_{1}$ nonzero elements and $\|{\mathbf{D}}_{1}{\mathbf{v}}\|_{2}\leq 1$, we have
$\sup_{\|{\mathbf{v}}\|_{2}=1}|{\mathbf{v}}^{\prime}{\mathbf{D}}_{1}{\mathbf{S}}_{R}{\mathbf{D}}_{1}{\mathbf{v}}|\leq\sup_{\|{\mathbf{v}}\|_{2}=1}\frac{{\mathbf{v}}^{\prime}{\mathbf{D}}_{1}}{\|{\mathbf{D}}_{1}{\mathbf{v}}\|_{2}}{\mathbf{S}}_{R}\frac{{\mathbf{D}}_{1}{\mathbf{v}}}{\|{\mathbf{D}}_{1}{\mathbf{v}}\|_{2}}\|{\mathbf{D}}_{1}{\mathbf{v}}\|_{2}^{2}\leq\sup_{{\mathbf{v}}\in S_{1}}|{\mathbf{v}}^{\prime}{\mathbf{S}}_{R}{\mathbf{v}}|,$
where the set $S_{1}$ is define in Lemma A.4. Then (IA.10) reduces to
$\sqrt{1-(\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1})^{2}}\leq\frac{2}{\lambda_{1}-\lambda_{2}}\sup_{{\mathbf{v}}\in S_{1}}|{\mathbf{v}}^{\prime}{\mathbf{S}}_{R}{\mathbf{v}}|.$
This completes the proof. $\Box$
The following lemma is from Theorem 2.8.1 of Vershynin (2018) or Lemma D.2 of Vu and Lei (2013).
Lemma A.5.
(Bernstein’s inequality). Let $Y_{1},...,Y_{N}$ be independent random variables with zero mean. Then
$P\left(\left|\sum_{i=1}^{N}Y_{i}\right|>t\right)\leq 2\exp\left(-c\min\left(\frac{t^{2}}{\sum_{i=1}^{N}\|Y_{i}\|_{\psi_{1}}^{2}},\frac{t}{\max_{i\leq N}\|Y_{i}\|_{\psi_{1}}}\right)\right),$
where
$\|Y_{i}\|_{\psi_{1}}=\inf_{K>0}\{K:E\exp(|Y_{i}|/K)\leq 2\}.$
Lemma A.6.
Let Assumptions 17 hold. For any ${\mathbf{v}}\in S_{1}$, we have
$P(|{\mathbf{v}}^{\prime}{\mathbf{e}}{\mathbf{e}}^{\prime}{\mathbf{v}}^{\prime}-E({\mathbf{v}}^{\prime}{\mathbf{e}}{\mathbf{e}}^{\prime}{\mathbf{v}}^{\prime})|>t)\leq 2\exp\left(-C\min\left(\frac{t^{2}}{N},t\right)\right).$
As a result,
$P(|{\mathbf{v}}^{\prime}{\mathbf{S}}_{e}{\mathbf{v}}^{\prime}-E({\mathbf{v}}^{\prime}{\mathbf{S}}_{e}{\mathbf{v}}^{\prime})|>t)\leq 2\exp\left(-C\min\left({NT^{2}t^{2}},NTt\right)\right).$
Proof. Without loss of generality, we assume the first $2s_{1}$ elements of ${\mathbf{v}}$ are nonzero, i.e. ${\mathbf{v}}=(v_{1},...,v_{2s_{1}},0,...,0)^{\prime}$. Then, ${\mathbf{v}}^{\prime}{\mathbf{e}}=({\mathbf{v}}^{\prime}\underline{{\mathbf{e}}}_{1},...,{\mathbf{v}}^{\prime}\underline{{\mathbf{e}}}_{N})$, where
${\mathbf{v}}^{\prime}\underline{{\mathbf{e}}}_{i}=\sum_{j=1}^{2s_{1}}v_{j}e_{i,j}.$
Define $Z_{i}={\mathbf{v}}^{\prime}\underline{{\mathbf{e}}}_{i}$, it follows from Assumption 7 that $Z_{i}$ is a sub-Gaussian random variable. By Lemma 2.7.6 of Vershynin (2018), $Z_{i}^{2}$ is a sub-exponential random variable and $\|Z_{i}^{2}\|_{\psi_{1}}\leq C$. Note that ${\mathbf{v}}^{\prime}{\mathbf{e}}{\mathbf{e}}^{\prime}{\mathbf{v}}=\sum_{i=1}^{N}Z_{i}^{2}$, by Lemma A.5, we have
$P\left(\left|\sum_{i=1}^{N}(Z_{i}^{2}-E(Z_{i}^{2}))\right|>t\right)\leq$ $2\exp\left(-c\min\left(\frac{t^{2}}{\sum_{i=1}^{N}\|Z_{i}^{2}\|_{\psi_{1}}^{2}},\frac{t}{\max_{i\leq N}\|Z_{i}^{2}\|_{\psi_{1}}}\right)\right),$
$\leq$ $2\exp\left(-C\min\left(\frac{t^{2}}{N},t\right)\right),$ (IA.11)
for some constant $C>0$. This completes the proof. $\Box$
Lemma A.7.
Let Assumptions 17 hold. For any vector ${\mathbf{v}}\in S_{1}$, we have
$P(|{\mathbf{v}}^{\prime}\boldsymbol{\Sigma}_{fe}{\mathbf{v}}|>x)\leq N\exp(-C(N\sqrt{T}x)^{\gamma})+\exp(-CNTx^{2})+\exp(-CNTx^{2}\exp(\frac{(N\sqrt{T}x)^{\gamma(1-\gamma)}}{(\log(N\sqrt{T}x))^{\gamma}})),$
where $\gamma=(1/\gamma_{1}+1)^{-1}$ for any $\gamma_{1}>0$.
Proof. Let $W_{j}={\mathbf{v}}^{\prime}\frac{{\mathbf{F}}}{\sqrt{T}}\boldsymbol{\lambda}_{j}$, by definition, we have
${\mathbf{v}}^{\prime}\boldsymbol{\Sigma}_{fe}{\mathbf{v}}=\frac{1}{NT}{\mathbf{v}}^{\prime}{\mathbf{F}}\boldsymbol{\Lambda}^{\prime}{\mathbf{e}}^{\prime}{\mathbf{v}}=\frac{1}{NT}\sum_{j=1}^{N}{\mathbf{v}}^{\prime}{\mathbf{F}}\boldsymbol{\lambda}_{j}\underline{{\mathbf{e}}}_{j}^{\prime}{\mathbf{v}}=\frac{1}{N\sqrt{T}}\sum_{j=1}^{N}W_{j}Z_{j},$
where $Z_{j}$ is defined as that in the proof of Lemma A.6. Note that $W_{j}$ and $Z_{j}$ are independent of each other and they all have finite variances. In addition, Assumption 5 implies that $Z_{j}$ are independent across $j=1,...,N$. By Assumption 7, $W_{j}Z_{j}$ satisfies condition (2.7) in Merlevède, Peligrad, and Rio (2011) with $\gamma_{2}=1$. Since $Z_{j}$ are independent over $j=1,...,N$, then condition (2.6) in Merlevède, Peligrad, and Rio (2011) is satisfied for any $\gamma_{1}>0$. By Theorem 1 of Merlevède, Peligrad, and Rio (2011), we have
$P(|\sum_{j=1}^{N}W_{j}Z_{j}|>x)\leq N\exp(-Cx^{\gamma})+\exp(-Cx^{2}/N)+\exp(-C\frac{x^{2}}{N}\exp(\frac{x^{\gamma(1-\gamma)}}{(\log(x))^{\gamma}})),$ (IA.12)
where $\gamma=(1/\gamma_{1}+1)^{-1}$. Then Lemma A.7 follows from (IA.12). $\Box$
Proof of Theorem 1. Note that
$\sup_{v\in S_{1}}|{\mathbf{v}}^{\prime}{\mathbf{S}}_{e}{\mathbf{v}}|\leq\sup_{v\in S_{1}}|{\mathbf{v}}^{\prime}({\mathbf{S}}_{e}-E{\mathbf{S}}_{e}){\mathbf{v}}|+\sup_{v\in S_{1}}|{\mathbf{v}}^{\prime}(E{\mathbf{S}}_{e}){\mathbf{v}}|=\Pi_{1}+\Pi_{2}.$ (IA.13)
By Assumption 5, we have $E{\mathbf{S}}_{e}=\sigma{\mathbf{I}}_{T}$, where
$\sigma=\frac{\sigma_{1}^{2}+...+\sigma_{N}^{2}}{NT}.$
As $\sigma\leq C/T$ for some constant $C>0$, we have
$\Pi_{2}=\sup_{v\in S_{1}}|{\mathbf{v}}^{\prime}(E{\mathbf{S}}_{e}){\mathbf{v}}|\leq C/T.$ (IA.14)
Now turn to $\Pi_{1}$. Note that $S_{1}=\{{\mathbf{v}}:\|{\mathbf{v}}\|_{2}=1\,\,\text{and}\,\,\|{\mathbf{v}}\|_{0}\leq 2s_{1}\}$. Let $\mathcal{S}_{2}^{T-1}$ be the set of $N$-dimensional unit vector and $\mathcal{B}_{0}^{d}$ be the set of vectors satisfiying $\|{\mathbf{v}}\|_{0}\leq d$. Then $S_{1}=\mathcal{S}_{2}^{T-1}\cap\mathcal{B}_{0}^{2s_{1}}$.
For every possible subset $I\subset\{1,...,T\}$ of size $2s_{1}$, it is well known that the minimal $\delta$-covering of $\mathcal{S}_{2}^{2s_{1}-1}$ in the Euclidean metric has cardinality at most $(1+2/\delta)^{2s_{1}}$. See, for example, Proposition D.2 in Vu and Lei (2013). Let $\mathcal{N}$ be the minimal $\delta$-covering set of $\mathcal{S}_{2}^{2s_{1}-1}$ with all possible subsets of $\{1,...,T\}$ with size $2s_{1}$, it is not hard to see that
$|\mathcal{N}|\leq\left(\begin{array}[]{c}N\\ 2s_{1}\end{array}\right)(1+\frac{2}{\delta})^{2s_{1}}\leq\left(\frac{eN}{2s_{1}}\right)^{2s_{1}}\left(1+\frac{2}{\delta}\right)^{2s_{1}},$
where we use the binomial coefficient bound for the first term. Note that for every ${\mathbf{x}}\in\mathcal{S}_{2}^{T-1}\cap\mathcal{B}_{0}^{2s_{1}}$, there exists ${\mathbf{y}}\in\mathcal{N}$ satisfying $\|{\mathbf{x}}-{\mathbf{y}}\|_{2}\leq\delta$ and ${\mathbf{x}}-{\mathbf{y}}\in\mathcal{B}_{0}^{2s_{1}}$. Therefore,
$\sup_{v\in S_{1}}|{\mathbf{v}}^{\prime}({\mathbf{S}}_{e}-E{\mathbf{S}}_{e}){\mathbf{v}}|\leq(1-2\delta)^{-1}\sup_{{\mathbf{v}}\in\mathcal{N}}|{\mathbf{v}}^{\prime}({\mathbf{S}}_{e}-E{\mathbf{S}}_{e}){\mathbf{v}}|.$
We choose $\delta=1/4$, by the Bonferroni-type inequality and Lemma A.6,
$P\left(\sup_{v\in S_{1}}|{\mathbf{v}}^{\prime}({\mathbf{S}}_{e}-E{\mathbf{S}}_{e}){\mathbf{v}}|>t\right)\leq$ $P\left((1-2\delta)^{-1}\sup_{{\mathbf{v}}\in\mathcal{N}}|{\mathbf{v}}^{\prime}({\mathbf{S}}_{e}-E{\mathbf{S}}_{e}){\mathbf{v}}|>t\right)$
$\leq$ $P\left(\sup_{{\mathbf{v}}\in\mathcal{N}}|{\mathbf{v}}^{\prime}({\mathbf{S}}_{e}-E{\mathbf{S}}_{e}){\mathbf{v}}|>t/2\right)$
$\leq$ $|\mathcal{N}|P\left(|{\mathbf{v}}^{\prime}({\mathbf{S}}_{e}-E{\mathbf{S}}_{e}){\mathbf{v}}|>t/2\right)$
$\leq$ $C\left(\frac{eN}{2s_{1}}\right)^{2s_{1}}9^{2s_{1}}\exp\left(-C\min\left({NT^{2}t^{2}},NTt\right)\right).$ (IA.15)
Therefore, we can obtain from (A) that
$\sup_{v\in S_{1}}|{\mathbf{v}}^{\prime}({\mathbf{S}}_{e}-E{\mathbf{S}}_{e}){\mathbf{v}}|\leq C\max\{\sqrt{\frac{s_{1}\log(N)}{NT^{2}}},\frac{s_{1}\log(N)}{NT}\}\leq C\sqrt{\frac{s_{1}\log(N)}{NT^{2}}}+C\frac{s_{1}\log(N)}{NT}.$
Combining (IA.14) and the above inequality, we have
$\sup_{v\in S_{1}}|{\mathbf{v}}^{\prime}{\mathbf{S}}_{e}{\mathbf{v}}|\leq C\left(\sqrt{\frac{s_{1}\log(N)}{NT^{2}}}+\frac{s_{1}\log(N)}{NT}+\frac{1}{T}\right).$ (IA.16)
By a similar argument and Lemma A.7, we can also show that
$\sup_{v\in S_{1}}|{\mathbf{v}}^{\prime}\boldsymbol{\Sigma}_{fe}{\mathbf{v}}|\leq C\left(\sqrt{\frac{s_{1}\log(N)}{NT}}\right).$ (IA.17)
By Assumption 4 that $\lambda_{1}-\lambda_{2}>c>0$, Theorem 1 follows from Lemma A.4, (IA.1), and the above two inequalities. This completes the proof. $\Box$
Proof of Theorem 2. According to Algorithms 1–3, we can repeat the proofs of Theorem 1 for the eigenvectors obtained in the subsequent iterations. So long as the previous eigenvector is consistent, we can similarly obtain the consistencies of the next few estimated eigenvectors. On the other hand, because the number of factors $r$ is finite, by (3), we have
$\rho(\widehat{\mathbf{V}}_{1},\widehat{\mathbf{V}}_{1})^{2}=2\mbox{tr}\left({\mathbf{I}}_{r}-\widehat{\mathbf{V}}^{\prime}{\mathbf{V}}{\mathbf{V}}^{\prime}\widehat{\mathbf{V}}\right).$
Note that the columns in the true factor matrix ${\mathbf{F}}$ are orthogonal, then the cross terms in $\widehat{\mathbf{V}}{\mathbf{V}}$ are of the same order as that in Theorem 1. We can apply the previous results to each component of the one on the right-hand side, and the result in Theorem 2 holds since $r$ is finite. See also the descriptions for different distances between two matrices in Equations (18)–(3) of the main text. This completes the proof. $\Box$
To show the consistency of the cross-validation method, we first introduce some notation. We denote ${\mathbf{F}}^{0}$ as the true factors with sparsity $s_{0}$, which is also the true one. ${\mathbf{F}}^{s}$ is the true factor matrix which keeps the largest $s$ elements in each column of ${\mathbf{F}}^{0}$ if $s\leq s_{0}$ and it is equal to ${\mathbf{F}}^{0}$ if $s\geq s_{0}$. $\widetilde{\mathbf{F}}_{1}^{s}$ is the estimated factor matrix with sparsity $s$ using the training sample. Define
$C_{NT}=\sqrt{\frac{s_{0}\log(N)}{NT}}+\frac{s_{0}\log(N)}{NT}+\frac{1}{T},$
which is the upper bound for $\|\widehat{\mathbf{F}}-{\mathbf{F}}\|_{F}/\sqrt{T}$ as shown in Theorem 2.
Proof of Theorem 3. Let $\widehat{\mathbf{F}}$ be the estimated sparse factors by the proposed algorithms. We apply the least-squares method to (3) and obtain
$\widehat{\boldsymbol{\Lambda}}^{\prime}=(\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}})^{-1}\widehat{\mathbf{F}}^{\prime}{\mathbf{X}}=(\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}})^{-1}\widehat{\mathbf{F}}^{\prime}{\mathbf{F}}\boldsymbol{\Lambda}^{\prime}+(\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}})^{-1}\widehat{\mathbf{F}}^{\prime}{\mathbf{e}}.$ (IA.18)
It follows from (IA.18) that
$\widehat{\boldsymbol{\lambda}}_{i}-{\mathbf{H}}_{s}\boldsymbol{\lambda}_{i}=(\frac{\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}}}{T})^{-1}\frac{\widehat{\mathbf{F}}^{\prime}{\mathbf{e}}_{i}}{T},$ (IA.19)
where ${\mathbf{H}}_{s}=(\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}})^{-1}\widehat{\mathbf{F}}^{\prime}{\mathbf{F}}$. By Assumption 7 and the results in Theorem 1, it is not hard to show that
$\max_{1\leq i\leq N}|\frac{\widehat{\mathbf{F}}^{\prime}{\mathbf{e}}_{i}}{T}|=O_{p}(\sqrt{\frac{s^{*}}{NT}}\log(N)+\frac{s^{*}\log(N)^{3/2}}{NT}+\sqrt{\frac{\log(N)}{T}}).$
If $s^{*}$ is finite, then it reduces to
$\max_{1\leq i\leq N}|\frac{\widehat{\mathbf{F}}^{\prime}{\mathbf{e}}_{i}}{T}|=O_{p}(\sqrt{\frac{\log(N)}{T}}).$
Then Theorem 3(i) holds.
Note that $\frac{\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}}}{T}\rightarrow_{p}{\mathbf{I}}_{r}$, by Assumptions 56 and the results in Theorem 2, we can show that
$\sqrt{T}(\widehat{\boldsymbol{\lambda}}_{i}-{\mathbf{H}}_{s}\boldsymbol{\lambda}_{i})\longrightarrow_{d}N(0,\sigma_{i}^{2}{\mathbf{I}}_{r}),$ (IA.20)
where $\sigma_{i}$ is defined in Assumption 5. This completes the proof. $\Box$
Lemma A.8.
Let Assumptions 17 hold. For any $s$ with $1\leq s\leq s_{0}$, we have
$R(s,\widetilde{\mathbf{F}}^{s})-R(s,{\mathbf{F}}^{s})=O_{p}(C_{N_{1}T}),$
where we ignore the subscript in (14) and denote $\widetilde{\mathbf{F}}^{s}$ as the estimated factor matrix from $\widetilde{\mathbf{X}}_{1}$ with sparsity $s$, and ${\mathbf{F}}^{s}$ as the true factor matrix retaining only keeps the top $s$ elements, both of which are normalized.
Proof. We only show it for $r=1$ since the proof is similar for $r>1$. Suppose ${\mathbf{F}}^{s}$ is the normalized true factor matrix such that ${\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{s}/T={\mathbf{I}}_{r}$. By (14), we have
$R(s,\widetilde{\mathbf{F}}^{s})-R(s,{\mathbf{F}}^{s})=\frac{1}{N_{2}T}\|\widetilde{\mathbf{X}}_{2}-\frac{1}{T}\widetilde{\mathbf{F}}^{s}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{X}}_{2}\|_{F}^{2}-\frac{1}{N_{2}T}\|\widetilde{\mathbf{X}}_{2}-\frac{1}{T}{\mathbf{F}}^{s}{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{X}}_{2}\|_{F}^{2}.$ (IA.21)
Note that
$\|\widetilde{\mathbf{X}}_{2}-\frac{1}{T}\widetilde{\mathbf{F}}^{s}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{X}}_{2}\|_{F}^{2}=\mbox{tr}\{(\widetilde{\mathbf{X}}_{2}-\frac{1}{T}\widetilde{\mathbf{F}}^{s}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{X}}_{2})(\widetilde{\mathbf{X}}_{2}-\frac{1}{T}\widetilde{\mathbf{F}}^{s}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{X}}_{2})^{\prime}\}$
and
$\|\widetilde{\mathbf{X}}_{2}-\frac{1}{T}{\mathbf{F}}^{s}{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{X}}_{2}\|_{F}^{2}=\mbox{tr}\{(\widetilde{\mathbf{X}}_{2}-\frac{1}{T}{\mathbf{F}}^{s}{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{X}}_{2})(\widetilde{\mathbf{X}}_{2}-\frac{1}{T}{\mathbf{F}}^{s}{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{X}}_{2})^{\prime}\}.$
By an elementary argument, we can show that
$R(s,\widetilde{\mathbf{F}}^{s})-R(s,{\mathbf{F}}^{s})=\frac{1}{N_{2}T}\mbox{tr}\{(\frac{1}{T}{\mathbf{F}}^{s}{\mathbf{F}}^{s}{{}^{\prime}}-\frac{1}{T}\widetilde{\mathbf{F}}^{s}\widetilde{\mathbf{F}}^{s}{{}^{\prime}})\widetilde{\mathbf{X}}_{2}\widetilde{\mathbf{X}}_{2}^{\prime}\}.$ (IA.22)
Since
$\|\frac{1}{T}{\mathbf{F}}^{s}{\mathbf{F}}^{s}{{}^{\prime}}-\frac{1}{T}\widetilde{\mathbf{F}}^{s}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}\|_{2}=\|\frac{1}{\sqrt{T}}({\mathbf{F}}^{s}-\widetilde{\mathbf{F}}^{s})\frac{{\mathbf{F}}^{s}{{}^{\prime}}}{\sqrt{T}}+\frac{\widetilde{\mathbf{F}}^{s}}{\sqrt{T}}\frac{({\mathbf{F}}^{s}-\widetilde{\mathbf{F}}^{s})^{\prime}}{\sqrt{T}}\|_{2}=O_{p}(C_{N_{1}T}),$ (IA.23)
and
$\|\widetilde{\mathbf{X}}_{2}\widetilde{\mathbf{X}}_{2}^{\prime}\|_{2}=O_{p}(N_{2}T),$
it follows from (IA.22) that
$R(s,\widetilde{\mathbf{F}}^{s})-R(s,{\mathbf{F}}^{s})\leq\frac{r}{N_{2}T}\|(\frac{1}{T}{\mathbf{F}}^{s}{\mathbf{F}}^{s}{{}^{\prime}}-\frac{1}{T}\widetilde{\mathbf{F}}^{s}\widetilde{\mathbf{F}}^{s}{{}^{\prime}})\widetilde{\mathbf{X}}_{2}\widetilde{\mathbf{X}}_{2}^{\prime}\|_{2}=O_{p}(C_{N_{1}T}).$
This completes the proof. $\Box$
Lemma A.9.
Let Assumptions 17 hold. For any $s$ with $s<s_{0}$, there exists a constant $\tau_{s}>0$ such that
$R(s,{\mathbf{F}}^{s})-R(s_{0},{\mathbf{F}}^{0})\geq\tau_{s},$
with probability tending to one as $N,T\rightarrow\infty$. ${\mathbf{F}}^{s}$ is the same as that in Lemma A.8 and ${\mathbf{F}}^{0}$ is the true factor matrix with true sparsity $s_{0}$.
Proof. By a similar argument as that in (IA.22), we can obtain that
$R(s,{\mathbf{F}}^{s})-R(s_{0},{\mathbf{F}}^{0})=\frac{1}{N_{2}T}\mbox{tr}\{[{\mathbf{F}}^{0}({\mathbf{F}}^{0}{{}^{\prime}}{\mathbf{F}}^{0})^{-1}{\mathbf{F}}^{0}{{}^{\prime}}-{\mathbf{F}}^{s}({\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{s})^{-1}{\mathbf{F}}^{s}{{}^{\prime}}]\widetilde{\mathbf{X}}_{2}\widetilde{\mathbf{X}}_{2}^{\prime}\},$ (IA.24)
where we may only impose one normalization condition on ${\mathbf{F}}^{s}$ or ${\mathbf{F}}^{0}$. For example, we may assume ${\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{s}/T={\mathbf{I}}_{r}$ and ${\mathbf{F}}^{0}={\mathbf{F}}^{s}+{\mathbf{G}}$, where $G$ is a nonzero vector when $s<s_{0}$. Note that
$\widetilde{\mathbf{X}}_{2}\widetilde{\mathbf{X}}_{2}^{\prime}={\mathbf{F}}^{0}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}+{\mathbf{F}}^{0}\boldsymbol{\Lambda}_{2}{{}^{\prime}}{\mathbf{E}}_{2}^{\prime}+{\mathbf{E}}_{2}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}+{\mathbf{E}}_{2}{\mathbf{E}}_{2}^{\prime},$ (IA.25)
where $\boldsymbol{\Lambda}_{2}$ and ${\mathbf{E}}_{2}$ are the loading matrix and idiosyncratic term of $\widetilde{\mathbf{X}}_{2}$. Then,
$\mbox{tr}\{[{\mathbf{F}}^{0}({\mathbf{F}}^{0}{{}^{\prime}}{\mathbf{F}}^{0})^{-1}{\mathbf{F}}^{0}{{}^{\prime}}-$ ${\mathbf{F}}^{s}({\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{s})^{-1}{\mathbf{F}}^{s}{{}^{\prime}}]\widetilde{\mathbf{X}}_{2}\widetilde{\mathbf{X}}_{2}^{\prime}\}$
$=$ $\mbox{tr}\{{\mathbf{F}}^{0}\boldsymbol{\Lambda}_{2}\boldsymbol{\Lambda}_{2}^{\prime}{\mathbf{F}}^{0}-{\mathbf{F}}^{s}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}-{\mathbf{F}}^{s}({\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{s})^{-1}{\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{G}}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}$
$+{\mathbf{F}}^{0}\boldsymbol{\Lambda}_{2}^{\prime}{\mathbf{E}}_{2}^{\prime}+{\mathbf{F}}^{0}({\mathbf{F}}^{0}{{}^{\prime}}{\mathbf{F}}^{0})^{-1}{\mathbf{F}}^{0}{{}^{\prime}}{\mathbf{E}}_{2}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}-{\mathbf{F}}^{s}({\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{s})^{-1}{\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{0}\boldsymbol{\Lambda}_{2}^{\prime}{\mathbf{E}}_{2}^{\prime}$
$-{\mathbf{F}}^{s}({\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{s})^{-1}{\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{E}}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}-{\mathbf{F}}^{s}({\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{s})^{-1}{\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{E}}_{2}{\mathbf{E}}_{2}^{\prime}\}$
$=$ $I+II,$ (IA.26)
where
$I=\mbox{tr}\{{\mathbf{F}}^{0}\boldsymbol{\Lambda}_{2}\boldsymbol{\Lambda}_{2}^{\prime}{\mathbf{F}}^{0}-{\mathbf{F}}^{s}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}-{\mathbf{F}}^{s}({\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{s})^{-1}{\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{G}}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}\},$
and $II$ is the remaining one. It is obvious that $II$ is of a smaller order than $I$, and we only consider $I$. Note that ${\mathbf{F}}^{0}={\mathbf{F}}^{s}+{\mathbf{G}}$, we have
$I=\mbox{tr}\{{\mathbf{G}}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{G}}^{\prime}-{\mathbf{F}}^{s}({\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{s})^{-1}{\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{G}}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{G}}^{\prime}\}.$
Since each column in ${\mathbf{F}}^{s}({\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{s})^{-1}{\mathbf{F}}^{s}{{}^{\prime}}$ is a hat matrix and ${\mathbf{G}}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{G}}^{\prime}$ is a positive semi-definite matrix, then there exists a constant $\tau_{s}>0$ such that $I\geq\tau_{s}N_{2}T$. Then Lemma A.9 follows from this result and (IA.24)–(A). This completes the proof. $\Box$
Lemma A.10.
Let Assumptions 17 hold. For any $s$ with $s\geq s_{0}$, we have
$R(s,\widetilde{\mathbf{F}}^{s})-R(s_{0},\widetilde{\mathbf{F}}^{0})=O_{p}(C_{N_{1}T}^{2}),$
where
Proof. First, we observe that
$|R(s,\widetilde{\mathbf{F}}^{s})-R(s_{0},\widetilde{\mathbf{F}}^{0})|\leq$ $|R(s,\widetilde{\mathbf{F}}^{s})-R(s_{0},{\mathbf{F}}^{0})|+|R(s_{0},{\mathbf{F}}^{0})-R(s_{0},\widetilde{\mathbf{F}}^{0})|$
$\leq$ $2\max_{s_{0}\leq s\leq smax}|R(s,\widetilde{\mathbf{F}}^{s})-R(s_{0},{\mathbf{F}}^{0})|.$ (IA.27)
Thus, it is sufficient to prove for each $s$ with $s\geq s_{0}$,
$|R(s,\widetilde{\mathbf{F}}^{s})-R(s_{0},{\mathbf{F}}^{0})|=O_{p}(C_{N_{1}T}^{2}).$ (IA.28)
Note that
$R(s,\widetilde{\mathbf{F}}^{s})-R(s_{0},{\mathbf{F}}^{0})=$ $\frac{1}{N_{2}T}\mbox{tr}\{[{\mathbf{F}}^{0}({\mathbf{F}}^{0}{{}^{\prime}}{\mathbf{F}}^{0})^{-1}{\mathbf{F}}^{0}{{}^{\prime}}-\widetilde{\mathbf{F}}^{s}(\widetilde{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{F}}^{s})^{-1}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}]\widetilde{\mathbf{X}}_{2}\widetilde{\mathbf{X}}_{2}^{\prime}\}$
$=$ $\frac{1}{N_{2}T}\mbox{tr}({\mathbf{L}}).$ (IA.29)
By (IA.25), we can show that
${\mathbf{L}}=$ ${\mathbf{F}}^{0}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}+{\mathbf{F}}^{0}\boldsymbol{\Lambda}_{2}^{\prime}{\mathbf{E}}_{2}^{\prime}+{\mathbf{F}}^{0}({\mathbf{F}}^{0}{{}^{\prime}}{\mathbf{F}}^{0})^{-1}{\mathbf{F}}^{0}{{}^{\prime}}{\mathbf{E}}_{2}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}+{\mathbf{F}}^{0}({\mathbf{F}}^{0}{{}^{\prime}}{\mathbf{F}}^{0})^{-1}{\mathbf{F}}^{0}{{}^{\prime}}{\mathbf{E}}_{2}{\mathbf{E}}_{2}^{\prime}$
$-\widetilde{\mathbf{F}}^{s}(\widetilde{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{F}}^{s})^{-1}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{0}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}-\widetilde{\mathbf{F}}^{s}(\widetilde{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{F}}^{s})^{-1}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{0}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}-\widetilde{\mathbf{F}}^{s}(\widetilde{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{F}}^{s})^{-1}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{E}}_{2}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}$
$-\widetilde{\mathbf{F}}^{s}(\widetilde{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{F}}^{s})^{-1}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{E}}_{2}{\mathbf{E}}_{2}^{\prime}.$ (IA.30)
We only show the upper bound for the dominant term since those for the rest are similar. Note that
$\mbox{tr}\{{\mathbf{F}}^{0}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}-$ $\widetilde{\mathbf{F}}^{s}(\widetilde{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{F}}^{s})^{-1}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{0}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}\}$
$=$ $\mbox{tr}\{{\mathbf{F}}^{0}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}-\widetilde{\mathbf{F}}^{s}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}-\widetilde{\mathbf{F}}^{s}(\widetilde{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{F}}^{s})^{-1}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s})\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}\}$
$=$ $\mbox{tr}\{({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s})\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}-\widetilde{\mathbf{F}}^{s}(\widetilde{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{F}}^{s})^{-1}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s})\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}\}$
$=$ $\mbox{tr}\{({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s})\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}-({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s})\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}$
$-\widetilde{\mathbf{F}}^{s}(\widetilde{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{F}}^{s})^{-1}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s})\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s}){{}^{\prime}}\}$
$=$ $\mbox{tr}\{({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s})\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s}){{}^{\prime}}-\widetilde{\mathbf{F}}^{s}(\widetilde{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{F}}^{s})^{-1}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s})\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s}){{}^{\prime}}\}$ (IA.31)
It is not hat to show that
$\|({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s})\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s}){{}^{\prime}}\|_{2}\leq N_{2}T\|\frac{({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s})}{\sqrt{T}}\frac{\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}}{N_{2}}\frac{({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s}){{}^{\prime}}}{\sqrt{T}}\|_{2}=O_{p}(N_{2}TC_{N_{1}T}^{2}).$
Then, we can show that
$\frac{1}{N_{2}T}\mbox{tr}\{{\mathbf{F}}^{0}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}-\widetilde{\mathbf{F}}^{s}(\widetilde{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{F}}^{s})^{-1}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{0}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}\}=O_{p}(C_{N_{1}T}^{2}).$ (IA.32)
By a similar argument for the remaining terms in (A), we can show that
$\frac{1}{N_{2}T}\mbox{tr}({\mathbf{L}})=O_{p}(C_{N_{1}T}^{2}),$
which implies (IA.28). This completes the proof. $\Box$
Proof of Theorem 4. We only consider the case when the number of partitions $J=1$ since it is similar for a genera $J>1$. We will prove that
$\lim_{N,T\rightarrow\infty}P(PC(s)<PC(s_{0}))=0,$
for all $s\neq s_{0}$ and $s\leq smax$. We first consider the case when $s<s_{0}$. Note that
$PC(k)-PC(s_{0})=R(s,\widetilde{\mathbf{F}}^{s})-R(s_{0},\widetilde{\mathbf{F}}^{0})-r(s_{0}-s)g(N_{1},T),$
and
$R(s,\widetilde{\mathbf{F}}^{s})-R(s_{0},\widetilde{\mathbf{F}}^{0})=[R(s,\widetilde{\mathbf{F}}^{s})-R(s,{\mathbf{F}}^{s})]+[R(s,{\mathbf{F}}^{s})-R(s_{0},{\mathbf{F}}^{0})]+[R(s_{0},{\mathbf{F}}^{0})-R(s_{0},\widetilde{\mathbf{F}}^{0})].$
It follows from Lemma A.8 that
$R(s,\widetilde{\mathbf{F}}^{s})-R(s,{\mathbf{F}}^{s})=O_{p}(C_{N_{1}T}),$
and
$R(s_{0},{\mathbf{F}}^{0})-R(s_{0},\widetilde{\mathbf{F}}^{0})=O_{p}(C_{N_{1}T}).$
Lemma A.9 implies that
$R(s,{\mathbf{F}}^{s})-R(s_{0},{\mathbf{F}}^{0})\geq\tau_{s}>0.$
Then,
$P(PC(s)-PC(s_{0})<0)=P(\tau_{s}+O_{p}(C_{N_{1}T})<r(s_{0}-s)g(N_{1},T))\rightarrow 0,$
since $g(N_{1},T)\rightarrow 0$ and $C_{N_{1}T}\rightarrow 0$.
Next, we consider the case when $s\geq s_{0}$. Similar,
$P(PC(s)-PC(s_{0})<0)=P(R(s_{0},\widetilde{\mathbf{F}}^{0})-R(s,\widetilde{\mathbf{F}}^{s})>r(s-s_{0})g(N_{1},T)).$
By Lemma A.10, we have
$R(s_{0},\widetilde{\mathbf{F}}^{0})-R(s,\widetilde{\mathbf{F}}^{s})=O_{p}(C_{N_{1}T}^{2}).$
Since $C_{N_{1}T}^{-1}g(N_{1},T)\rightarrow\infty$, we have that
$P(R(s_{0},\widetilde{\mathbf{F}}^{0})-R(s,\widetilde{\mathbf{F}}^{s})>r(s-s_{0})g(N_{1},T))\rightarrow 0,$
as $N,T\rightarrow\infty$. This completes the proof. $\Box$
B Description of the Identified Sparse Factors
In this section, we report the identified significant factors of stock returns over the time horizon from January 1, 2004, to December 31, 2016, studied in the empirical analysis. A comprehensive list of dates with significant systematic risk factors, the reasons for these factors, and their associated time points are provided in Table LABEL:Table-ft.
Table IA.I: Significant sparse factors over the time horizon from January 1, 2004, to December 31, 2016. Dates are formatted as yyyymmdd. The signs in the Factor Return column indicate whether the significant time factor resulted in a positive ($+$) or negative (--) return on that day. Reasons in the Reason column are extracted from daily reports on CNN Money (www.money.cnn.com) after the market closes each trading day. The Time Factor column summarizes the corresponding factor for each trading day with significant common risk.
Date Factor Return Reason Time Factor
20070227 -- A big decline in Chinese stocks, weakness in some key readings on the US economy and news that Vice President Dick Cheney was the apparent target in a Taliban suicide bombing attack in Afghanistan China, Economic indicators, Global events (terrorist attack)
20070803 -- Credit market fears, sparked by Wall Street bank Bear Stearns Credit risk, Market sentiment
20070918 $+$ After the Federal Reserve cut a key short-term interest rate by a half-percentage point Government policies
20071107 -- Renewed credit market fears Credit risk, Market sentiment
20071128 $+$ Expectations the Federal Reserve will continue its rate cutting campaign and on strength from the embattled financials Government policies, Market sentiment
20071211 -- After the Federal Reserve cut the fed funds rate by a quarter-percentage point, as expected, but disappointed some investors looking for a bigger cut Government policies, Market sentiment
20080104 -- After a weaker-than-expected December jobs report exacerbated worries that the economy may be falling into recession Economic indicators, Market sentiment
20080117 -- Recession worries following comments from Federal Reserve Chairman Ben Bernanke, Merrill Lynch’s big quarterly loss and weak readings on the housing and manufacturing sectors Market sentiment, Company-specific factors, Economic indicators
20080123 $+$ The Federal Reserve stepped in Tuesday and announced an emergency intermeeting interest rate cut, a decision that initially had a mixed impact on stocks, but helped pave the way for Wednesday’s bounce back Government policies
20080205 -- After a report showing a big slowdown in the services sector of the economy and cautionary comments from a Fed official amplified fears that a recession is underway or imminent Company-specific factors, Government policies
20080229 -- After AIG’s record loss added to worries about the financial sector and more weak economic news intensified fears about a recession Company-specific factors
20080311 $+$ Announcement: the Federal Reserve will lend up to \$$200$ billion to banks and lenders as a means of loosening up tight credit markets Government policies
20080318 $+$ After the Federal Reserve cut the fed funds rate by three-quarters of a percentage point, surprising investors looking for a larger cut Government policies
20080401 $+$ Investors cheered signs that the companies hit hardest by the credit market crisis seem to be working through the problems Market sentiment
20080606 -- After oil prices spiked more than \$currency-dollar\\$\$$11$ a barrel and the May jobs report showed a big jump in the unemployment rate Oil price, Economic indicators
20080626 -- Selling accelerated following a record surge in oil prices Oil price
20080708 $+$ Falling oil prices and a stronger dollar Oil price
20080709 -- More worries about Freddie Mac and Fannie Mae’s ability to raise capital exacerbated credit market and corporate profit jitters Market sentiment
20080716 $+$ Encouraging news from the banking and airline sectors Falling oil prices also helped spark a strong stock market rally Economic indicators, Oil price
20080717 $+$ A string of large oil price declines and more encouraging earnings results from the financial sector Oil price, Economic indicators
20080724 -- Renewed fears about the battered housing market and rising unemployment Market sentiment
20080729 $+$ Strong financial reports, rising consumer confidence and falling oil prices Economic indicators, Oil price
20080805 $+$ Oil prices fell sharply and investors appeared to take solace in the Federal Reserve’s assessment of the nation’s economy Oil price, Government policies
20080904 -- Mixed retail sales, lower oil prices and dour labor market readings amplified worries about a global economic slowdown Economic indicators, Oil price
20080909 -- Worries about Lehman Brothers’ ability to raise capital, and about the extent of AIG’s mortgage-related losses, exacerbated broad recession fears Company-specific factors
20080915 -- After one of the most calamitous days in US financial services history resulted in Bank of America’s \$$50$ billion deal to buy Merrill Lynch and the bankruptcy filing of Lehman Brothers Company-specific factors
20080917 -- The government’s emergency rescue of AIG amplified fears about the stability of financial markets Government policies
20080918 $+$ On a CNBC report that the government is working on a more permanent solution to absorbing bad debt Government policies
20080919 $+$ The government’s plan to help rescue banks from toxic mortgage debt soothed investors at the end of a gut-churning week on Wall Street Government policies
20080922 -- As investors worried about the specifics of the government’s \$$700$ billion bailout plan and rocketing oil prices - which saw its biggest one-day dollar gain ever Market sentiment
20080929 -- After the House rejected the government’s \$$700$ billion bank bailout plan Government policies
20080930 $+$ On bets that Congress will pass a version of the government’s \$$700$ billion package, following Monday’s crushing defeat Market sentiment, Government policies
20081002 -- As frozen credit markets and weak economic reports amplified jitters ahead of the House vote on the \$$700$ billion bank rescue plan Market sentiment, Government policies
20081006 -- As the \$$700$ billion bank bailout plan and European government attempts to prop up faltering banks failed to comfort panicky investors Market sentiment, Government policies
20081007 -- As the Federal Reserve’s plan to loosen credit markets failed to temper investor pessimism Market sentiment, Government policies
20081009 -- As panicked investors dumped stocks across the board Market sentiment
20081013 $+$ As investors bet that the worst of the credit crisis is over, following a series of global initiatives announced over the last few days Market sentiment
20081015 -- A weak retail sales report and dour forecasts from the Federal Reserve, coupled with sober comments from Fed Chairman Ben Bernanke, sent stocks tumbling Economic indicators, Government policies
20081016 $+$ As the lowest oil prices in more than a year gave investors a reason to scoop up shares battered in the recent market selloff Oil price
20081020 $+$ As investors welcomed talk of a second economic stimulus plan and an improvement in key lending rates Market sentiment, Government policies
20081021 -- As mixed corporate earnings reports gave investors a reason to retreat after the previous session’s big rally Economic indicators
20081022 -- As weak earnings and slumping oil prices amplified fears of a global recession Oil price
20081024 -- As Wall Street joined a worldwide market slump on bets that a recession is imminent - if not already under way Market sentiment
20081027 -- As recession jitters outweighed relief that the government’s programs to shore up the financial system have gotten underway Market sentiment
20081028 $+$ As investors dove back into stocks near the end of one of the worst months in Wall Street history Market sentiment
20081031 $+$ Capping off a strong week at the end of one of the worst months in Wall Street history Market sentiment
20081104 $+$ As millions of Americans battered by the weakened economy turned out to vote for the next President of the United States Market sentiment
20081105 -- As Barack Obama’s historic victory gave way to renewed worries about the struggling economy Market sentiment
20081106 -- As fears of a prolonged recession sent investors running for the exits Market sentiment
20081110 -- As ongoing recession fears overshadowed any relief about China’s \$$586$ billion stimulus plan and the government’s revamping of its deal with AIG China, Government policies
20081111 -- As recession fears trumped a new government and mortgage industry plan to help troubled homeowners Market sentiment, Government policies
20081112 -- As investors bet that a long and deep recession is on the horizon Market sentiment
20081113 $+$ As the major stock gauges bounced back from levels not seen since 2003 Market sentiment
20081114 -- As the worst retail sales on record ignited fears of a long recession Economic indicators
20081117 -- As investors eyed Citigroup’s massive job losses and a weak manufacturing report, while awaiting the fate of a potential bailout for the automakers Company-specific factors, Economic indicators
20081119 -- As ongoing anxiety about the economy and uncertainty about the future of the auto industry weighed on the market Market sentiment
20081120 -- As fears of a prolonged recession sparked a massive selloff Market sentiment
20081121 $+$ After reports surfaced that President-elect Barack Obama will nominate New York Federal Bank President Timothy Geithner as his new Treasury secretary Government policies
20081124 $+$ As Citigroup’s massive rescue package and President-elect Obama’s picks for his economic team pushed investors off the sidelines Company-specific factors, Government policies
20081126 $+$ As investors scooped up stocks hit in the recent selloff, ahead of the Thanksgiving holiday Market sentiment
20081201 -- As investors bailed out following confirmation that the US is mired in a recession and indications that it’s likely to continue for some time Market sentiment
20081202 $+$ As investors welcomed signs that the automakers might get a bailout after all Government policies
20081204 -- As a rash of job cuts at major companies added to jitters ahead of the November jobs report Economic indicators
20081205 $+$ After a brutal November employment report, as investors extended the recent trend of buying despite the bad news Economic indicators
20081208 $+$ As investors welcomed President-elect Barack Obama’s plan to create jobs and revive the economy, and reports that government help for the automakers is on the way Government policies
20081211 -- On worries that the \$currency-dollar\\$\$$14$ billion auto rescue bill won’t pass in the Senate due to Republican opposition Government policies
20081216 $+$ After the Federal Reserve cut a key short-term interest rate to the lowest level on record, and signaled it had more tools available to help the economy as the recession stretches on Government policies, Market sentiment
$2$- Amid concerns about fourth-quarter corporate earnings, falling oil prices and ongoing woes in the auto industry Market sentiment, Oil price, Economic indicators
20081230 $+$ As investors scooped up a variety of shares hit hard in the 2008 stock market battering Market sentiment
20090102 $+$ With investors starting off a new year on the right foot, after an abysmal 2008 Market sentiment
20090106 $+$ As investors looked beyond the Federal Reserve’s dour outlook on the economy and instead scooped up shares hit in last year’s big selloff Market sentiment
20090107 -- After weak labor market reports and dour forecasts from Alcoa and Intel gave investors reasons to retreat after the recent rally Economic indicators, Company-specific factors
20090109 -- After a government report showed another big monthly drop in payrolls, resulting in the biggest annual job loss since just after World War II Economic indicators
20090112 -- By concerns about Citigroup’s potential deal with Morgan Stanley - and the start of the fourth-quarter earnings reporting period Company-specific factors
20090114 -- As a bleak retail sales report and more dour news from the banking sector amplified fears of a prolonged recession Economic indicators
20090120 -- As investors looked beyond President Barack Obama’s historic inauguration to the battered economy he inherits Market sentiment
20090121 $+$ As investors welcomed IBM’s earnings and scooped up bank shares hit hard in the recent retreat Company-specific factors
20090122 -- As a management shakeup at Bank of America and Microsoft’s earnings disappointment weighed on investor sentiment Company-specific factors
20090128 $+$ As investors took comfort in reports that the Obama administration and the Federal Reserve are taking steps to get credit flowing again and help staunch the economic slowdown Government policies
20090129 -- Following more dire news on earnings, housing and employment Economic indicators
20090130 -- As investors eyed abysmal reports on economic growth and quarterly earnings Economic indicators
20090206 $+$ As optimism about the government’s economic stimulus bill and the new version of the bank bailout plan countered unease following the brutal January jobs report Government policies, Economic indicators
20090210 -- As the government’s bank rescue plan failed to reassure investors burned by the 14-month old recession Government policies
20090217 -- On fears that the government’s efforts to slow the recession won’t be sufficient Market sentiment
20090219 -- As fears of a prolonged recession sent stock investors heading for the exits Market sentiment
20090223 -- As investors continue to worry that the government’s efforts to slow the recession won’t be sufficient Market sentiment
20090224 $+$ After comments from Fed Chairman Ben Bernanke that downplayed bank takeover fears helped to spark a big rally Government policies, Market sentiment
20090227 -- On worries about the government taking a bigger chunk of Citigroup and a bleak reading on the economy, again touching 12-year lows Government policies, Economic indicators
20090302 -- After insurance company American International Group’s huge quarterly loss added to worries about the financial sector and the economy Company-specific factors
20090304 $+$ Following reports that China’s economy may be improving and as government officials unveiled details of the \$$75$ billion foreclosure fix China
20090305 -- As investors waded through more grim news: GM said its survival is in doubt, bank shares took a beating, and Citigroup fell below a buck Company-specific factors
20090310 $+$ After Citigroup cooled some worries about its future and regulators said they may reinstate a key trading rule Company-specific factors, Government policies
20090312 $+$ As investors scooped up banks and other shares hit in a selloff that left the Dow at 12-year lows Market sentiment
20090317 $+$ As investors continued to dig out from 12-year lows Market sentiment
20090318 $+$ After the Federal Reserve said it would buy up to \$300 billion in long-term government bonds Government policies
20090320 -- As investors pulled back after the recent run Market sentiment
20090323 $+$ After Treasury’s plan to buy up billions in bad bank assets and a better-than-expected existing home sales report raised hopes that the economy is stabilizing Government policies, Economic indicators
20090326 $+$ As the March market run shows its legs Market sentiment
20090330 -- As auto and bank woes spark a selloff after the rally Economic indicators
20090402 $+$ After key accounting rule that has impact on banks is changed G-20 also in focus Government policies
20090407 -- After a four-week advance, on worries about banks and autos and the start of the quarterly reporting period Market sentiment
20090409 $+$ After Wells Fargo forecast a nearly \$$3$ billion quarterly profit, adding to hopes that the banking sector is stabilizing Company-specific factors
20090414 -- After a weaker-than-expected retail sales report gave investors a reason to retreat following a five-week run Company-specific factors, Market sentiment
20090416 $+$ The major stock gauges touch the best levels in months on JPMorgan Chase earnings Company-specific factors
20090420 -- On worries about financial sector earnings, despite Bank of America’s better-than-expected quarterly results Market sentiment
20090421 $+$ As worries about corporate results were countered by renewed hopes that the financial sector is closer to stabilizing Market sentiment
20090424 $+$ After Ford, Microsoft and American Express reported results that met or topped analysts’ expectations Company-specific factors
20090429 $+$ After the Federal Reserve held interest rates steady, as expected, but issued a slightly more upbeat economic outlook Government policies
20090504 $+$ As a better-than-expected housing market report intensified hopes that the economy is closer to stabilizing Economic indicators
20090508 $+$ After a government report showed employers cut fewer jobs than expected last month Economic indicators
20090511 -- As investors took a step back after propelling the major stock gauges by more than $30\%$ each in just two months Market sentiment
20090513 -- After a weaker-than-expected retail sales report gave investors a reason to retreat Economic indicators
20090518 $+$ After positive news about the US housing market, including an upbeat profit forecast from Lowes, as well as an upgrade of Bank of America Economic indicators, Company-specific factors
20090526 $+$ After a report showing consumer confidence hit an eight-month high offset dismal housing news Market sentiment
20090601 $+$ As better-than-expected readings on manufacturing activity raised hopes that a global economic recovery is brewing Economic indicators
20090615 -- As weaker oil prices and more geopolitical unrest raised worries that the recession may not be waning as soon as some had hoped Oil price, Global events (Geopolitical events)
20090622 -- As the World Bank’s weak outlook on global growth and a selloff in commodity prices sent investors heading for the exits Market sentiment
20090702 -- After a worse-than-expected jobs report hammered hopes that the economy is close to stabilizing Economic indicators
20090713 $+$ As investors welcomed an analysts’ improved outlook on Goldman Sachs one day ahead of its quarterly report Market sentiment, Company-specific factors
20090715 $+$ After Intel’s forecast for a second-half pickup and the Federal Reserve’s improved outlook reassured wary investors Company-specific factors
20090723 $+$ As investors welcomed better-than-expected quarterly results and home sales Economic indicators
20090817 -- As worries that nervous consumers will pressure a fragile recovery dragged stocks lower after a five-month advance Market sentiment
20090901 -- On worries that the market gains have raced ahead of any economic recovery Market sentiment
20091001 -- After a bigger-than-expected rise in weekly jobless claims and a weaker-than-expected reading on manufacturing sparked worries about the pace of the economic recovery Economic indicators
20091028 -- As a weaker-than-expected new home sales report added to questions about the strength of the economic recovery Economic indicators
20091029 $+$ As a strong report on economic growth in the third quarter reassured investors that the recovery is on track Economic indicators
20091030 -- As investors dumped a variety of shares at the end of a rough week and choppy month on Wall Street Market sentiment
20100204 -- A growing debt crisis in Europe Europe (Debt crisis)
20100427 -- After Standard & Poors cut Greece’s debt rating to junk and lowered Portugal’s debt rating, Europe (Debt crisis)
20100504 -- On worries that the global recovery could suffer if Europe’s efforts to contain Greece’s debt problems don’t succeed, and if China’s efforts to slow its booming economy go too far Europe (Debt crisis), China
20100506 -- Fears about the spread of the European debt crisis Europe (Debt crisis)
20100510 $+$ After European officials approved a nearly $1$ trillion rescue plan to contain the debt crisis in troubled nations and stabilize the euro Government policies
20100520 -- Worries about how the European debt crisis and slump in the euro will impact the global recovery fueled the selling Europe (Debt crisis)
20100527 $+$ After Chinese officials dismissed reports that they’re reviewing their nation’s investment in European bonds amid concerns about the continent’s debt problems China
20100601 -- As worries about the global economic outlook overshadowed better-than-expected readings on the US economy Market sentiment
20100604 -- After a government report showed employers added fewer jobs than expected last month and the euro plunged to a new 4-year low, reviving worries about the health of the European economyEconomic Indicators Economic indicators, Europe (Debt crisis)
20100610 $+$ As concerns over Europe’s debt crisis and its impact on the global recovery were calmed by a sharp boost in Chinese exports and a strengthening euro Europe (Debt crisis), China
20100629 -- After a big drop in consumer confidence and signs of a bigger slowdown in the global economy Market sentiment
20100707 $+$ As investors came back after the recent bloodletting, spurred on by State Street’s improved earnings forecast Economic indicators
20100716 -- After financial firms Bank of America and Citigroup reported weaker quarterly revenue and a plunge in consumer sentiment revived concerns about the economic outlook Company-specific factors
20100722 $+$ After better-than-expected earnings and forecasts from 3M, Caterpillar, AT&T and UPS helped reassure investors about the pace of the economic recovery Economic indicators, Company-specific factors
20100811 -- After a report showed the US trade gap widened, and foreign data cast doubt on overseas demand for American goods Economic indicators
20100901 $+$ As investors cheered signs of strength in the manufacturing sector Economic indicators
$2$- Libya’s escalating political crisis sparked a sharp sell-off in US stocks, as oil prices continued to skyrocket Global events (geopolitical events)
20110601 -- Weak economic data has started to snowball Economic indicators
20110802 -- As fears about a weak US economy were enflamed after investors got another disappointing economic report - this time on consumer spending Economic indicators
20110804 -- As fear about the global economy spooked investors Market sentiment
20110808 -- As the debt crisis in Europe, lackluster economic news and a downgrade to the US credit rating Europe (Debt crisis), Credit risk
20110809 $+$ After the Federal Reserve said it will keep interest rates exceptionally low until 2013 Government policies
20110810 -- As investors were confronted with mounting fears about Europe’s ongoing debt crisis, this time in France Europe (Debt crisis)
20110811 $+$ On positive earnings and labor market news Economic indicators
20110818 -- As renewed concerns about the US and global economies Market sentiment
20110823 $+$ Following a report from the FDIC that showed the number of failing banks shrank for the first time in nearly five years Economic indicators
20110829 $+$ A Greek bank deal, a solid US consumer spending report and relief that Hurricane Irene caused less damage than expected Government policies
20110902 -- After a government report showing no job growth in August stoked fears that the US may be headed into another recession Economic indicators
20110907 $+$ The gains came as concerns over Europe’s debt crisis eased and investors geared up for President Obama’s highly anticipated jobs speech Thursday evening Europe (Debt crisis)
20110909 -- As bad news out of Europe kept piling up Europe (Debt crisis)
20110921 -- Fed disappoints Government policies
20110922 -- Fear factor Market sentiment
20110928 -- Worries over stalled global growth prospects and concerns that European leaders may not be moving fast enough to solve the region’s debt problems Market sentiment
20110930 -- As investors remain worried about the debt crisis in Europe and the outlook for global economic growth Market sentiment
20111003 -- With worries about Greece’s solvency still in the spotlight Market sentiment
20111004 $+$ Bear market bounce Market sentiment
20111010 $+$ As investors cheered a pledge from European leaders to unveil a plan for solving the eurozone’s debt crisis by the end of the month Government policies
20111017 -- As worries about Europe’s debt crisis dominated Market sentiment, Europe (Debt crisis)
20111018 $+$ Following a report suggesting that Europe’s bailout fund may get a big boost Market sentiment, Government policies
20111027 $+$ After European Union leaders agreed to expand Europe’s bailout fund and take major losses on Greek bonds Government policies
20111031 -- As investors continued to scrutinize the eurozone debt deal Europe (Debt crisis)
20111101 -- New fears about the fate of the European rescue plan Market sentiment
20111109 -- Italy fears Europe (Debt crisis)
20111123 -- As eurozone fears rumbled on and a preliminary report showed that Chinese manufacturing slowed sharply Europe (Debt crisis), China
20111128 $+$ On robust Black Friday sales Economic indicators
20111130 $+$ After the Federal Reserve said it will work with other central banks to support the global economy Government policies
20111220 $+$ As concerns about the European debt crisis eased and investors welcomed signs of strength in the US housing market Market sentiment, Economic indicators
20120601 -- As Europe’s debt crisis remains unresolved and the US economy is showing new signs of distress Europe (Debt crisis), Market sentiment
20120629 $+$ A deal among European leaders to help struggling eurozone banks buoyed global markets Government policies
20130415 -- Following the news of explosions at the Boston Marathon Global events (terrorist attack)
20150824 -- Deep fears about China’s economic slowdown China
20150826 $+$ The huge rally represents a rebound following six days of dramatic selling that was driven by serious concerns about how China’s slowing economy will impact the rest of the world China
20150901 -- Following more fears about a slowdown in China’s economy China
20160113 -- Low oil prices, as concerns about global economic slowdown Oil price
20160129 $+$ A realization that if the US avoids a recession – as most economists think it will – beaten-down stocks could be a good buy Market sentiment
20160624 -- British voters chose to leave the European Union Europe (Brexit)
20160627 -- British voters chose to leave the European Union Europe (Brexit)